### v1

# Implicit Bias in Matrix Factorization and its Explicit Realization in a New Architecture

Abstract
Gradient descent for matrix factorization is known to exhibit an implicit bias toward approximately low-rank solutions. While existing theories often assume the boundedness of iterates, empirically the bias persists even with unbounded sequences. We thus hypothesize that implicit bias is driven by divergent dynamics markedly different from the convergent dynamics for data fitting. Using this perspective, we introduce a new factorization model: $X\approx UDV^{\top}$, where $U$ and $V$ are constrained within norm balls, while $D$ is a diagonal factor allowing the model to span the entire search space. Our experiments reveal that this model exhibits a strong implicit bias regardless of initialization and step size, yielding truly (rather than approximately) low-rank solutions. Furthermore, drawing parallels between matrix factorization and neural networks, we propose a novel neural network model featuring constrained layers and diagonal components. This model achieves strong performance across various regression and classification tasks while finding low-rank solutions, resulting in efficient and lightweight networks.
1 Introduction
The Burer–Monteiro (BM) factorization (Burer & Monteiro, 2003) is a classical technique for obtaining low-rank solutions in optimization. It can be viewed as a simple neural network that uses a single layer of hidden neurons under linear activation. Indeed, say one has the factorization $X=UV^{T}$ where $U\in\mathbb{R}^{d\times r}$ and $V\in\mathbb{R}^{c\times r}$, then $U$ and $V$ can be thought of as the weights of the first and second layers, while $r$ represents the number of hidden neurons. However, despite the similarity suggested by this view, there is a clear distinction between BM factorization and neural networks about how the rank $r$ is chosen. In BM factorization, $r$ is typically chosen to be small, close to the rank of the desired solution. Neural networks, on the other hand, often succeed even in overparametrized settings where $r$ is large.
Recent findings of implicit regularization in matrix factorization narrows the gap between these two perspectives. For instance, in an effort to explain the empirical success of overparametrized neural networks, Gunasekar et al. (2017) demonstrate that gradient descent (with certain parameter selection) on BM factorization tends to converge toward approximately low-rank solutions even when $r=d$. Based on this observation, they conjecture that “with small enough step sizes and initialization close enough to the origin, gradient descent on full-dimensional factorization converges to the minimum nuclear norm solution.”
In a follow-up work, however, Razin & Cohen (2020) presented a counter-example demonstrating that implicit regularization in BM factorization cannot be explained by minimal nuclear norm, or in fact any norm. Specifically, they showed that there are instances where the gradient method applied to BM factorization yields a diverging sequence, and all norms thus grow toward infinity. Intriguingly, despite this divergence, they found that the rank of the estimate decreases toward its minimum.
Although this phenomenon might seem surprising initially, it is not uncommon for diverging sequences to follow a structured path. A prime example is the Power Method, the fundamental algorithm for finding the largest eigenvalue and eigenvector pair of a matrix. Starting from a random initial point $x_{0}$, the Power Method iteratively updates the estimate by multiplying it with the matrix. This process amplifies the component of the vector that aligns with the direction of the dominant eigenvector more than the other components, progressively leading $x_{k}$ to align with this eigenvector. In practical implementations, $x_{k}$ is scaled after each iteration to avoid numerical issues from divergence.
This perspective serves as the foundational motivation guiding our approach. Specifically, our key insight is that the implicit regularization in BM factorization (and neural networks) is driven by divergent dynamical behavior. This is markedly different from the standard (convergent) optimization dynamics helping with the data fitting. In this context, we hypothesize that these forces do not merely coexist but actively compete, influencing model behavior and performance in fundamentally conflicting ways. Our main goal in the development of this paper is to devise an approach that unravels these competing forces.
1.1 Overview of main contributions
  • $\blacksquare$
    A novel formulation for matrix factorization. We model $X=UDV^{\top}$, where $U$ and $V$ are constrained within Frobenius norm balls. Projection onto this ball results in a scaling step similar to the Power Method. The middle term $D$ is a diagonal matrix that allows the model to explore the entire search space despite $U$ and $V$ being bounded.
    Empirically we demonstrate that the gradient method applied to the proposed formulation exhibits a pronounced implicit bias toward low-rank solutions. We compare our formulation against standard BM factorization with two unconstrained factors. Specifically, we investigate key factors such as step size and initialization, which prior work suggests might be contributing to implicit bias. We find that our factorization approach largely obviates the need to rely on these conditions: it consistently finds truly (rather than approximately) low-rank solutions across a wide range of initializations and step-sizes in our experiments. We believe these findings should be of broader interest to research on implicit bias.
  • $\blacksquare$
    A novel neural network architecture. Motivated by the strong bias for low-rank solutions of the proposed factorization, we subsequently extend it to deep neural networks. Specifically, we do so by adding constrained layers and diagonal components. We demonstrate numerically that this constrained model performs on par with, or even better than, the standard architecture across various regression and classification tasks. Importantly, even here, our approach exhibits bias towards low-rank solutions, resulting in a natural pruning procedure to obtain compact, lightweight networks without compromising performance.
2 Related Work
Burer-Monteiro factorization. BM factorization is initially proposed for solving semi-definite programs (Burer & Monteiro, 2003, 2005; Boumal et al., 2016, 2020; Cifuentes, 2021) and is recognized for its efficiency in addressing low-rank matrix optimization problems, see (Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017, 2018; Hsieh et al., 2018; Sahin et al., 2019; Lee et al., 2022; Yalçın et al., 2023) and the references therein. Building on the connections between training problems for (non-linear) two-layer neural networks and convex (copositive) formulations (see (Pilanci & Ergen, 2020; Ergen & Pilanci, 2020; Sahiner et al., 2021) and the references therein), Sahiner et al. (2023) recently introduced BM factorization to solve convex formulations for various neural network architectures, including fully connected networks with ReLU and gated ReLU (Fiat et al., 2019) activations, two-layer convolutional neural networks (LeCun et al., 1989a), and self-attention mechanisms based on transformers (Vaswani et al., 2017).
Implicit regularization. While neural network-based systems are rapidly emerging as a dominant technology, the mechanisms underlying their generalization capabilities are yet to be fully understood. One promising line of research aims to explain this success through the concept of ‘implicit regularization,’ which is induced by the optimization methods and formulations used during neural network training (Neyshabur et al., 2014, 2017; Neyshabur, 2017). Several studies have explored matrix factorization to investigate implicit bias in linear neural networks (Gunasekar et al., 2017; Arora et al., 2018; Razin & Cohen, 2020; Belabbas, 2020; Li et al., 2021). Le & Jegelka (2022) extended these results to the final linear layers of nonlinear ReLU-activated feedforward networks with fully connected layers and skip connections. More recently Timor et al. (2023) investigated implicit regularization in ReLU networks. Much of the existing work focuses on gradient flow dynamics in the limit of infinitesimal learning rates. In particular, Gidel et al. (2019) examined discrete gradient dynamics in two-layer linear neural networks, showing that the dynamics progressively learn solutions of reduced-rank regression with a gradually increasing rank.
Constrained neural networks. Regularizers are frequently used in neural network training to prevent overfitting and improve generalization (Goodfellow et al., 2016), or to achieve structural benefits such as sparse and compact network architectures (Scardapane et al., 2017). However, it is conventional to apply these regularizers as penalty functions in the objective rather than as hard constraints, addressing them within gradient-based optimization via their (sub)gradients. This approach is likely favored due to the ease of implementation, as pre-built functions are readily available in common neural network packages. Several independent studies have also applied proximal methods across different networks and applications, finding that proximal-based methods tend to yield solutions with more pronounced structures (Bai et al., 2019; Yang et al., 2020; Yun et al., 2021; Yurtsever et al., 2021; Yang et al., 2022). Structural regularization in the form of hard constraints, however, appears to be rare in neural network training. One notable exception is in the context of neural network training with the Frank-Wolfe algorithm (Pokutta et al., 2020; Zimmer et al., 2022; Macdonald et al., 2022), where constraints are necessary due to the algorithm’s requirement for a bounded domain.
There are other use cases of constraints in neural networks. For instance, constraints can be applied to ensure adherence to real-world conditions in applications where they are necessary (Pathak et al., 2015; Márquez-Neila et al., 2017; Jia et al., 2017; Kervadec et al., 2019; Weber et al., 2021), or to incorporate physical laws in physics-informed neural networks (Raissi et al., 2019; Lu et al., 2021; Patel et al., 2022). In these cases, the feasible set is typically complex and difficult to project onto; therefore, optimization algorithms like primal-dual methods or augmented Lagrangian techniques are used. Constraints are also present in lifted neural networks, a framework where the training problem is reformulated in a higher-dimensional ‘lifted’ space. In this space, conventional activation functions like ReLU, used in the original formulation, are expressed as constraints (Askari et al., 2018; Sahiner et al., 2021; Bartan & Pilanci, 2021; Prakhya et al., 2024).
Pruning. Neural networks are often trained in an over-parameterized regime to enhance generalization and avoid getting stuck in poor local minima. However, these models can suffer from excessive memory and computational resource demands, making them less efficient for deployment in real-world applications (Chang et al., 2021). Several compression techniques have been studied in the literature, including parameter quantization (Krishnamoorthi, 2018), knowledge distillation (Gou et al., 2021), and pruning.
Pruning reduces the number of parameters, resulting in more compact and efficient models that are easier to deploy. The literature on pruning is extensive, with diverse methods proposed with distinct characteristics. Various criteria are used to determine which weights to prune, including second-order derivative-based methods (LeCun et al., 1989b; Hassibi & Stork, 1992), magnitude-based pruning (Janowsky, 1989; Han et al., 2015), saliency heuristics (Mozer & Smolensky, 1988; Lee et al., 2018), and matrix or tensor factorization-based techniques (Xue et al., 2013; Sainath et al., 2013; Jaderberg et al., 2014; Lebedev et al., 2015; Swaminathan et al., 2020), among others. A comprehensive review of pruning methods is beyond the scope of this paper due to space limitations and the diversity of approaches. For more detailed reviews, we refer to (Reed, 1993; Blalock et al., 2020; Cheng et al., 2024), and the references therein.
3 Matrix Factorization with a Diagonal Component
Consider the matrix sensing problem, where the goal is to recover a positive semidefinite (PSD) matrix $X\in\mathbb{S}_{+}^{d\times d}$ from a set of linear measurements $\smash{b=\mathcal{A}(X)\in\mathbb{R}^{n}}$. We define $\smash{\mathcal{A}:\mathbb{R}^{d\times d}\rightarrow\mathbb{R}^{n}}$ through symmetric measurement matrices $\smash{A_{1},\ldots,A_{n}\in\mathbb{S}^{d\times d}}$, such that $\mathcal{A}(X)=[\langle A_{1},X\rangle~{}\cdots~{}\langle A_{n},X\rangle]^{\top}$ and $\mathcal{A}^{\top}y=\sum_{i=1}^{n}y_{i}A_{i}.$ We particularly focus on the data-scarce setting where $n\ll d^{2}$. A popular example is the matrix completion problem, where the goal is to reconstruct a matrix $X$ from a subset of its entries. This problem is inherently under-determined; however, successful recovery is possible if $X$ is known to be low-rank (Candes & Recht, 2012).
Remark 1.
We focus on the recovery problem of a PSD matrix for simplicity; note that a non-square matrix sensing problem can also be reformulated as a PSD matrix sensing problem through a simple transformation (Park et al., 2017).
The problem described above can be cast as the following rank-constrained optimization problem:
$\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad f(X):=\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{subj.\ to}\quad\mathrm{rank}(X)\leq r.$ (1)
Although rank-constrained matrix optimization problems are typically NP-hard, various methods have been developed to provide practical approximations. Examples include hard thresholding algorithms (Jain et al., 2010; Goldfarb & Ma, 2011; Kyrillidis & Cevher, 2014), convex relaxation techniques (Candes & Recht, 2012; Recht et al., 2010), and BM factorization (Burer & Monteiro, 2003; Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017).
The main idea behind BM factorization is to reparametrize the decision variable $X$ as $UU^{\top}$, where the factor $U\in\mathbb{R}^{d\times r}$, and $r$ is a positive integer that controls the rank of the resulting product. Problem (1) can then be reformulated as:
$\min_{U\in\mathbb{R}^{d\times r}}\quad\frac{1}{2}\|\mathcal{A}(UU^{\top})-b\|_{2}^{2}.$ (2)
Despite the fact that finding the global minimum of (2) remains challenging, a local solution can be approximated using the gradient method (Lee et al., 2016). Initializing at $U_{0}\in\mathbb{R}^{d\times r}$, perform the iteration:
$U_{k+1}=U_{k}-\eta\nabla_{U}f(U_{k}U_{k}^{\top}),$ (3)
where $\eta>0$ is the step-size, and the gradient $\nabla_{U}f$ is computed as
$\nabla_{U}f(UU^{\top})=2\nabla f(UU^{\top})U=2\mathcal{A}^{\top}(\mathcal{A}(UU^{\top})-b)U.$
Selecting the factorization rank $r$ is a critical decision that might impact the solution. A small $r$ may lead to spurious local minima, resulting in inaccurate outcomes (Waldspurger & Waters, 2020). Conversely, a large $r$ might weaken rank regularization, rendering the problem underdetermined. Conventional wisdom in BM factorization suggests finding a moderate choice that balances these two extremes. However, a key observation in (Gunasekar et al., 2017) is that the gradient method applied to (2) exhibits a tendency towards approximately low-rank solutions (i.e., a decaying singular value spectrum) even when $r=d$. Below, we restate their conjecture:
Conjecture in (Gunasekar et al., 2017). Suppose gradient flow (i.e., gradient descent with an infinitesimally small step-size) is initialized at a full-rank matrix arbitrarily close to the origin. If the limit of the gradient flow, $X_{\text{GF}}=UU^{\top}$, exists and is a global optimum of (1) with $\mathcal{A}(X_{\text{GF}})=b$, then $X_{\text{GF}}$ is the minimal nuclear-norm solution to (1).
Remark 2.
This conjecture outlines three conditions for implicit bias: a small step-size, near-origin full-rank initialization, and consistent measurements ($\mathcal{A},b$). However, these conditions are often unconventional. For example, small step-sizes are typically avoided because they slow down convergence. Similarly, near-origin initialization is counterintuitive, as the origin is a trivial spurious stationary point of (2).
3.1 The Proposed Factorization
We propose reparameterizing $X=UDU^{\top}$, where $U\in\mathbb{R}^{d\times r}$ is constrained to have a bounded norm, and $D\in\mathbb{R}^{r\times r}$ is a non-negative diagonal matrix:
$\min_{\begin{subarray}{c}U\in\mathbb{R}^{d\times r}\\ D\in\mathbb{R}^{r\times r}\end{subarray}}~{}\frac{1}{2}\|\mathcal{A}(UDU^{\top})-b\|_{2}^{2}\quad\mathrm{s.t.}\quad\|U\|_{F}\leq\alpha,~{}~{}D_{ii}\geq 0,~{}~{}D_{ij}=0,~{}~{}\forall i~{}\text{and}~{}\forall j\neq i,$ (4)
where $\alpha>0$ is a model parameter. When the problem is well-scaled, for instance through basic preprocessing with data normalization, we found that $\alpha=1$ is a reasonable choice.
Placing in multiple factors and with constraints, we perform projected-gradient updates on $U$ and $D$ with step-size $\eta>0$:
$U_{k+1}$ $=\Pi_{U}\left(U_{k}-\eta\nabla_{U}f(U_{k}D_{k}U_{k}^{\top})\right)$ (5)
$D_{k+1}$ $=\Pi_{D}\left(D_{k}-\eta\nabla_{D}f(U_{k}D_{k}U_{k}^{\top})\right),$
where $\Pi_{U}$ and $\Pi_{D}$ are projections for the constraints in (4); while the gradients are
$\nabla_{U}f(UDU^{\top})$ $=2\nabla f(UDU^{\top})UD,$
$\nabla_{D}f(UDU^{\top})$ $=U^{\top}\nabla f(UDU^{\top})U.$
3.2 Numerical Experiments on Matrix Factorization
We present numerical experiments comparing the empirical performance of the proposed approach with the classical BM factorization. Specifically, we examine the impact of initialization and step-size on the singular value spectrum of the resulting solution.
We set up a synthetic matrix completion problem to recover a PSD matrix $\smash{X_{\natural}=U_{\natural}U_{\natural}^{\top}}$, where the entries of $\smash{U_{\natural}\in\mathbb{R}^{100\times 3}}$ are independently and identically (iid) drawn from the standard Gaussian distribution. We randomly sample $n=900$ entries of $X_{\natural}$ and store them in the vector $b\in\mathbb{R}^{n}$. The goal is to recover $X_{\natural}$ from $b$ by solving problems (2) and (4). For initialization, we generate $U_{0}\in\mathbb{R}^{d\times d}$ with entries drawn iid from standard Gaussian distribution and rescale it to have the Frobenius norm $\xi>0$ (we investigate the impact of $\xi$). We initiate $D_{0}$ from the identity matrix.
The results are shown in Figure 1. First, we examine the impact of step-size. To this end, we fix $\xi=10^{-2}$ and test different values of $\eta$. In the left panel, we plot the objective residual as a function of iterations. As expected, we observe that a smaller step-size slows down convergence. In the right panel, we plot the singular value spectrum of the results attained after $10^{6}$ iterations. We observe no direct connection between step-size and implicit bias in BM factorization.
Impact of step-size ($\eta$), in noiseless setting, with fixed initialization.
Impact of initial distance to origin $(\xi)$, in noiseless setting, with fixed step-size.
Figure 1: Impact of step-size and initialization on implicit bias. Solid lines represent our UDU factorization, while dashed lines denote the classical BM factorization. [Left] Objective residual vs. iterations. [Right] Singular value spectrum after $10^{6}$ iterations. In all cases, UDU produces truly low-rank solutions, whereas the classical approach results in approximate low-rank structures.
Next, we investigate the impact of initialization. We fix the step-size at $\eta=10^{-2}$ and evaluate the effect of varying $\xi$. We observe a correlation between the implicit bias of the BM factorization and $\xi$, which determines the initial distance from the origin. Initializing closer to the origin in the classical BM factorization yields solutions with a faster spectral decay. Notably, the UDU factorization demonstrates a strong implicit bias toward truly low-rank solutions, regardless of the choice of $\eta$ or $\xi$.
We provide additional experiments on matrix factorization problems in the Appendices. Specifically, Section A.1 considers the matrix completion problem with noisy measurements, where $b$ is perturbed with Gaussian noise. The results remain consistent with the noiseless case: the UDU model exhibits an implicit bias toward truly low-rank solutions, while the classical BM factorization yields approximately low-rank solutions, with spectral decay strongly influenced by initialization.
In Section A.2, we present the evolution of the factors $U$ and $D$ over the iterations. Interestingly, the UDU model produces rank-revealing solutions, with the columns of $U$ converging to zero in certain directions. Specifically, $U$ tends to grow along certain directions, while the rescaling induced by projection onto the bounded constraint shrinks other coordinates. This behavior aligns with the mechanism of the power method and offers insights into its connection with divergent forces.
Additionally, we present numerical experiments on a matrix sensing problem arising in phase retrieval image recovery in Section A.3. As before, the UDU framework consistently promotes low-rank solutions, and this structural bias significantly enhances the quality of the recovered image, as demonstrated by our results.
4 Feedforward Neural Networks with Diagonal Hidden Layers
This section extends our approach to neural networks. Consider a dataset comprising $n$ data points $(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathbb{R}^{d}\times\mathbb{R}^{c}$. We first define a three-layer neural network defined as
$\phi(\mathbf{x}):=\sum_{j=1}^{m}\mathbf{v}_{j}w_{j}\mathbf{u}_{j}^{\top}\mathbf{x}~{}~{}\approx~{}~{}\mathbf{y}.$ (6)
The first and third layers are fully connected, and the middle is a diagonal layer, as illustrated in Figure 2. Drawing parallels between our matrix factorization model in (4) and neural network training, we impose Euclidean norm constraints on the weights of the fully connected layers. Under these conditions, the training problem can be formulated as follows:
$\min_{\begin{subarray}{c}\mathbf{u}_{j}\in\mathbb{R}^{d}\\ \mathbf{v}_{j}\in\mathbb{R}^{c}\\ w_{j}\in\mathbb{R}\end{subarray}}$ $\frac{1}{2n}\sum_{i=1}^{n}\|\sum_{j=1}^{m}\mathbf{v}_{j}\,w_{j}\,\mathbf{u}_{j}^{\top}\mathbf{x}_{i}-\mathbf{y}_{i}\|_{2}^{2}$ (7)
$\mathrm{subj.to}$ $\sum_{j=1}^{m}\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~{}~{}\sum_{j=1}^{m}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1,~{}~{}\text{and }~{}~{}w_{j}\geq 0;\quad\text{for all}~{}~{}j=1,\ldots,m.$
The norm constraints in our training problem can be interpreted as a stronger form of weight decay, one of the most commonly used regularization techniques in neural networks, which lends further justification to our formulation. We refer to this neural network structure as UDV.
Figure 2: UDV structure. The weights in diagonal layer $D$ are denoted as $w_{j}$.
4.1 Numerical Experiments on Neural Networks
In this section, we test the proposed UDV framework on regression and classification tasks, comparing it with fully connected two-layer neural networks (denoted as UV in the subsequent text) using both linear and ReLU activation functions. This comparison is fair in terms of computational cost, as the cost incurred by the diagonal layer —which can also be viewed as a parameterized linear activation function— is negligible. We demonstrate a strong empirical bias toward low-rank solutions in all our experiments. We also present a proof-of-concept use case of this strong bias, combined with an SVD-based pruning strategy, to produce compact networks.
4.1.1 Implementation Details
Computing environment. All classification tasks were conducted on an NVIDIA A100 GPU with four cores of the AMD Epyc 7742 processor, while regression tasks were conducted on a single core of an Intel Xeon Gold 6132 processor. We used Python 3.9.5 and PyTorch 2.0.1.
Datasets. We used two datasets for the regression tasks: House Prices - Advanced Regression Techniques (HPART) (Anna Montoya, 2016) and New York City Taxi Trip Duration (NYCTTD) (Risdal, 2017). We allocated 80% of the data for the training and reserved the remaining 20% for validation. Following (Huang, 2003), we set the number of hidden neurons in the diagonal layer $\smash{m=\texttt{round}\big{(}\sqrt{(c+2)d}+2\sqrt{d/(c+2)}\big{)}}$. This results in a network structure ($d$-$m$-$c$) of 79-26-1 for HPART, and 12-10-1 for NYCTTD.
For classification tasks, we used the normalized MNIST dataset (LeCun et al., 2010). We applied transfer learning by replacing the classifier layers of three advanced neural networks –MaxViT-T (Tu et al., 2022), EfficientNet-B0 (Tan & Le, 2019), and RegNetX-32GF (Radosavovic et al., 2020)– with UDV, while using pre-trained weights from ImageNet-1K (Deng et al., 2009). Specifically, we retained all layers up to the first fully connected layer of the classifier and replaced the subsequent layers. The number of hidden neurons in the diagonal layer was set to as $m=\texttt{floor}(\tfrac{2}{3}d)$. This results in a UDV network structure ($d$-$m$-$c$) of 512-341-10 for MaxViT-T, 1280-853-10 for EfficientNet-B0, and 2520-1680-10 for RegNetX-32GF.
Loss function. We used mean squared error for regression and cross-entropy loss for classification.
Optimization methods. We tested the results using four different optimization algorithms for training: Adam (Kingma & Ba, 2014), Mini-Batch Gradient Descent (MBGD) (LeCun et al., 2002), NAdam (Dozat, 2016), and Mini-Batch Gradient Descent with Momentum (MBGDM) (Sutskever et al., 2013). For classification, we used different batch sizes for different models: 128 for MaxViT-T and RegNetX-32GF, and 384 for EfficientNet-B0.
We tuned the step size for all models and optimization algorithms: For regression tasks, we tested step sizes $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$, $2$, $3$. Larger step sizes ($1$, $2$, $3$) were often excluded for the UV model due to divergence. In classification experiments, we tested learning rates $10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$ with Adam and NAdam; and we tested $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$, $2$, $3$, $5$ with MBGD and MBGDM.
Training Procedure. UV and UDV models were initialized identically, with the diagonal elements of $D$ initialized using Kaiming Uniform Initialization, consistent with the default initialization for fully connected layers in PyTorch. The models were trained for 200 epochs on HPART, 50 epochs on NYCTTD, and 70 epochs on MNIST; and results were averaged over 1000 random seeds for HPART, 100 for NYCTTD, and 1 for MNIST to ensure robustness. The validation loss, used as a generalization metric in regression, was averaged over the final 20 epochs for the HPART dataset and the final 5 epochs for the NYCTTD dataset. Similarly, validation accuracy for classification tasks was averaged over the last 5 epochs to ensure stability in the reported values.
Repository. The code is available at https://github.com/Titanium-H/UDV
4.1.2 Low-rank Bias in Neural Network Training
Table 1 presents the validation loss (for regression) or validation accuracy (for classification) of the UDV model compared to the classical UV model with linear and ReLU activation functions. For each configuration (dataset and model architecture), the results are obtained by selecting the best algorithm and learning rate pair. Moreover, Figure 3 illustrates the singular value spectrum of the solutions corresponding to each entry in these tables. We focus on the singular values from the $U$ and $UD$ layers, as they generate the primary data representation, while omitting the $V$ layer, which serves as the feature selection layer and is a tall matrix by definition, given that $c\ll m$ in most cases. Collectively, these results show that the UDV framework achieves competitive prediction accuracy while exhibiting a strong implicit bias toward low-rank solutions, as indicated by the faster decay in the singular value spectrum.
Table 1: Model performance using different models. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset HPART NYCTTD MNIST
UDV
$1.304\times 10^{-3}$
Adam: $10^{-3}$
$5.248\times 10^{-6}$
NAdam: $10^{-4}$
M: 99.67%
MBGDM: $10^{-2}$
E: 99.63%
MBGDM: $10^{-1}$
R: 99.74%
MBGDM: $10^{-2}$
UV
$1.333\times 10^{-3}$
Adam: $10^{-3}$
$5.251\times 10^{-6}$
Adam: $10^{-3}$
M: 99.69%
MBGDM: $10^{-2}$
E: 99.60%
Adam: $10^{-3}$
R: 99.66%
MBGD: $10^{0}$
UV-ReLU
$1.167\times 10^{-3}$
Adam: $10^{-3}$
$5.323\times 10^{-6}$
NAdam: $10^{-3}$
M: 99.68%
NAdam: $10^{-4}$
E: 99.68%
MBGDM: $10^{-1}$
R: 99.73%
MBGD: $10^{0}$
Figure 3: Singular value spectrum corresponding to the solutions reported in Table 1.
4.1.3 Reducing Network Size with SVD-based Pruning
Efficient and lightweight feed-forward layers are crucial for real-world applications. For instance, the Apple Intelligence Foundation Models (Gunter et al., 2024) recently reported that pruning hidden dimensions in feed-forward layers yields the most significant gains in their foundation models. Building on this insight, we leverage the inherent low-rank bias of the UDV architecture through an SVD-based pruning strategy to produce compact networks without sacrificing performance.
A low-rank solution was observed when applying SVD to $UD$ layers:
$UD=\texttt{USV}^{\top},\quad\texttt{U}\in\mathbb{R}^{d\times m},\quad\texttt{S}\in\mathbb{R}^{m\times m},\quad\texttt{V}^{\top}\in\mathbb{R}^{m\times m}.$ (8)
By dropping small singular values in S, these matrices can be truncated to $\smash{\bar{\texttt{U}}\in\mathbb{R}^{d\times r}}$, $\smash{\bar{\texttt{S}}\in\mathbb{R}^{r\times r}}$ and $\bar{\texttt{V}}^{\top}\in\mathbb{R}^{r\times m}$, where $0<r<m$. Consequently, ($m-r$) neurons can be pruned, and new weight matrices are assigned:
$\bar{U}=\bar{\texttt{U}}\in\mathbb{R}^{d\times r},\quad\bar{D}=\bar{\texttt{S}}\in\mathbb{R}^{r\times r},\quad\bar{V}=\bar{\texttt{V}}^{T}V\in\mathbb{R}^{r\times c}.$ (9)
We applied this pruning strategy on the models from Table 1. The left part of Figure 4 presents an example comparing the generalization capability of pruned models. For comparison, we created compact models by training from scratch with a reduced number of neurons $m$ in the hidden layer. The performance change for these models is shown in the right panel of Figure 4. Although our pruned networks derived from UDV demonstrate that models with significantly fewer parameters can still achieve strong generalization, these compressed architectures are often more challenging to optimize directly within the reduced space, consistent with prior findings in the literature (Arora et al., 2018; Chang et al., 2021). We omit the results for retraining with the UV model, as they show similar trends to UDV in this context, though UDV generally exhibits superior generalization.
Figure 4: Comparison of SVD-based pruning and re-training compact networks on the HPART dataset using the NAdam algorithm with a learning rate of $10^{-3}$. Negative percentages indicate improvements over the baseline. SVD-based pruning demonstrates that the UDV leads to a compact model without performance degradation, while retraining shows that the UDV achieves better generalization in a compact model.
4.1.4 Further Details and Discussions
Our findings in experiments on neural networks align with the results observed in matrix factorization. A key distinction, however, was the use of different optimization algorithms, including stochastic gradients and momentum steps, in neural network experiments. Despite these differences, the UDV architecture consistently demonstrated a strong bias toward low-rank solutions. Additional experiments and details can be found in the Appendices, with key results summarized below:
  • In the early stages of this research, we explored four variants of UDV, each differing slightly in their constraints. We selected the version presented in (7), as it generally exhibits the most pronounced decay in the singular value spectrum. For completeness, details of the other three variants are provided in Section B.1.
  • Section B.2 provides additional results for the experiment described in Section 4.1.2. Specifically, we present results analogous to those in Table 1 and Figure 3, but focusing exclusively on the MBGDM algorithm. These results exhibit similar trends, reinforcing consistency across different methods. Additionally, comprehensive performance comparisons across all algorithms and models are provided in Tables SM2, SM3, SM4 and SM5 in the Appendices.
  • Additional results on SVD-based pruning are provided in Section B.3. We demonstrate that the UDV framework consistently achieves low-rank solutions across various problem configurations. Furthermore, we analyze the effect of learning rate on the singular value spectrum, similar to the analysis in Figure 1, but applied to neural network experiments. This analysis confirms that the UDV framework produces low-rank solutions across a broad range of learning rates.
  • Section B.4 extends the UDV framework by incorporating ReLU activation. Preliminary experiments with the UDV-ReLU model indicate that it also tends to yield low-rank solutions, similar to those observed in the original UDV framework.
  • Prior work on implicit bias in neural networks suggests that increasing depth enhances the tendency toward low-rank solutions (Arora et al., 2019; Feng et al., 2022), raising the question of whether the pronounced bias in the UDV framework is just a consequence of adding a diagonal layer. To investigate this, we conducted experiments comparing the UDV model to fully connected three-layer networks, as detailed in Section B.5. Additionally, we included a UDV model without constraints in these comparisons. The results indicate that this bias cannot be attributed solely to depth, highlighting the critical role of explicit constraints.
  • Section B.6 compares the spectral decay of the UDV network to that induced by classical weight decay regularization in two- and three-layer networks. The results show that while weight decay promotes singular value decay, it can not reproduce the strong decay observed in the UDV model.
5 Conclusions
We proposed a new matrix factorization framework, inspired by the observation that implicit bias is driven by dynamics—potentially divergent—that are distinct from those leading to convergence of the objective function. This framework constrains the factors within Euclidean norm balls and introduces a middle diagonal factor to ensure the search space is not restricted. Numerical experiments demonstrate that this approach significantly strengthens the low-rank bias in the solution.
To explore the broader applicability of our findings, we designed an analogous neural network architecture with three layers, constraining the fully connected layers and adding a diagonal hidden layer, referred to as UDV. Extensive experiments show that the proposed UDV architecture achieves competitive performance compared to standard fully connected networks, while inducing a structured solution with a strong bias toward low-rank representations. Additionally, we explored the utility of this low-rank structure by applying an SVD-based pruning strategy, illustrating how it can be leveraged to construct compact networks that are more efficient for downstream tasks.
Acknowledgments
Alp Yurtsever and Yikun Hou were supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Part of computations were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre. Another part of computations were conducted using the resources of High Performance Computing Center North (HPC2N) which is mainly funded by The Kempe Foundations and the Knut and Alice Wallenberg Foundation. Suvrit Sra acknowledges generous support from the Alexander von Humboldt (AvH) foundation.
Appendix A Additional Details on Matrix Factorization Experiments
A.1 Matrix Completion under Gaussian Noise
We conducted similar experiments to those in Section 3.2 also using noisy measurements: Let $b^{\natural}=\mathcal{A}(X^{\natural})$ represent the true measurements, and assume $b=b^{\natural}+\omega$, where $\omega\in\mathbb{R}^{n}$ is zero-mean Gaussian noise with a standard deviation of $\sigma=10^{-2}\|b\|_{2}$. The results, shown in Figure SM1, remain consistent with the noiseless case. $UDU^{\top}$ exhibits implicit bias toward truly low-rank solutions, while the classical BM factorization yields approximately low-rank solutions, with the spectral decay influenced by initialization.
Impact of step-size $(\eta)$, under Gaussian noise, with fixed initialization.
Impact of initial distance to origin $(\xi)$, under Gaussian noise, with fixed step-size.
Figure SM1: Impact of step-size and initialization on implicit bias. Solid lines represent our UDU factorization, while dashed lines denote the classical BM factorization. [Left] Objective residual vs. iterations. [Right] Singular value spectrum after $10^{6}$ iterations.
A.2 UDU Factorization Produces Rank-Revealing Solutions
The proposed factorization method naturally produces rank-revealing solutions. The columns of $U$ converge to zero in certain directions, resulting in truly low-rank solutions. Specifically, $U$ tends to grow along specific directions while the rescaling induced by projection onto the bounded constraint shrinks other coordinates. This behavior resembles the mechanism of the power method and provides insights into its connection with divergent forces.
In Figure SM2, we illustrate the evolution of the column norms of $U$ from the matrix completion experiment described in Section 3.2. Figure SM3 displays the corresponding entries in the diagonal factor $D$. For comparison, Figure SM4 shows the column norms of $U$ obtained from the same experiment using the standard BM factorization. The results are shown for $\xi=10^{-1}$ and $\eta=10^{-1}$ over $10^{5}$ iterations, but the findings are similar across other parameter settings. We also repeat the experiment with the noisy measurements described in Section A.1. The results are shown in Figures SM5, SM6 and SM7.
Figure SM2: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM3: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization. The x-axis represents indices.
Figure SM4: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization, shown for comparison. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM5: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM6: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents indices.
Figure SM7: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization with noisy measurements, shown for comparison. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
A.3 Phase Retrieval Image Recovery
We conduct a numerical experiment on the matrix sensing problem arising in phase retrieval image recovery (Candes et al., 2013). The objective is to recover a signal from its quadratic measurements of the form $y_{i}=|\langle a_{i},x\rangle|^{2}$. Although the standard maximum likelihood estimators lead to a non-convex optimization problem due to the quadratic terms, the problem can be reformulated as minimizing a convex function under a rank constraint in a lifted space. By denoting the lifted variable as $X=xx^{\top}$, the measurements can be expressed as
$\smash{y_{i}=\langle a_{i}^{\top}x,a_{i}^{\top}x\rangle=\langle a_{i}a_{i}^{\top},xx^{\top}\rangle:=\langle A_{i},X\rangle,}$
which allows us to reformulate the problem as
$\smash{\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{subj.\ to}\quad\mathrm{rank}(X)\leq 1,}$
representing a special case of problem (1).
In this experiment, we use a gray-scale image of size $16\times 16$ pixels, selected from the Pixel Art dataset (Elgazar, 2024), as the signal $x\in\mathbb{R}^{n}$ to recover, where $n=256$. We generate a synthetic measurement system by sampling $a_{1},\ldots,a_{m}$ from the standard Gaussian distribution, with $m=2n$. We then solve problems (2) and (4). We initialized $U_{0}\in R^{n\times n}$ with entries drawn iid from the standard Gaussian distribution, then rescaled to have a unit Frobenius norm. For $D_{0}\in R^{n\times n}$, we used the identity matrix. We used step-size $\eta=\frac{1}{L}$ where $L$ denotes the smoothness constant. After solving the problem, we recover $x\in\mathbb{R}^{n}$ from the lifted variable $X\in\mathbb{R}^{n\times n}$ by selecting its dominant eigenvector.
Figure SM8 demonstrates that the proposed method consistently identifies low-rank solutions, in line with the results of our other experiments. Figure SM9 displays the recovered image, demonstrating that the proposed method achieves higher quality than BM factorization within the same number of iterations, which can be attributed to the low-rank structure of the solution.
Figure SM8: Comparison of singular value spectrum in recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
Figure SM9: Comparison of recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
Appendix B Additional Details on Neural Network Experiments
B.1 Design Variants of the UDV Architecture
When designing our network architecture, we considered four variants, including the one presented in Section 4. The three additional models were defined by the following sets of constraints:
$\sum_{j=1}^{m}\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~{}~{}\sum_{j=1}^{m}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1$ (UDV-s)
$\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~{}~{}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1~{}~{}~{}w_{j}\geq 0;\quad\text{for all}~{}~{}j=1,\ldots,m$ (UDV-v1)
$\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~{}~{}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1;\quad\text{for all}~{}~{}j=1,\ldots,m$ (UDV-v2)
In detail, UDV-s is identical to UDV but omits the non-negativity constraints on the diagonal layer. UDV-v1, on the other hand, enforces row/column-wise norm constraints instead of the Frobenius norm used in UDV, while retaining the non-negativity constraints on the diagonal elements. Finally, UDV-v2 is identical to UDV-v1 but without the non-negativity constraints on the diagonal layer.
We only report results from UDV, as defined in equation (7), in the main text. However, we provide results for the other variants in the Appendices for completeness and to illustrate the impact of different constraint settings on model performance.
Figure SM11 and Figure SM12 show that UDV consistently finds a low-rank solution. Generally, UDV exhibits the most pronounced decaying pattern in singular values. Additionally, we extended our experiments to include full-batch training on the HPART and NYCTTD datasets. Although full-batch training converges more slowly than stochastic (mini-batch) methods, it exhibits a similar singular value decay pattern, confirming our earlier observations.
B.2 Comparison of Model Performance with the MBGDM Algorithm
In Section 4.1, we presented the performance of each model (UDV, UV, and UV-ReLU) in terms of generalization power (measured by validation loss or validation accuracy) along with the corresponding singular value spectrum, as shown in Table 1 and Figure 3. The results were obtained by selecting the optimal algorithm and learning rate pair for each model. Here, we conduct the same experiment, but focus exclusively on the MBGDM algorithm. The results, presented in Table SM1 and Figure SM10 are similar to the ones in Section 4.1, showing similar trends, highlighting consistency across methods.
Table SM1: Model performance using MBGDM. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset HPART NYCTTD MNIST
UDV
$1.345\times 10^{-3}$
MBGDM: $10^{-1}$
$5.248\times 10^{-6}$
MBGDM: $10^{-1}$
M: 99.67%
MBGDM: $10^{-2}$
E: 99.63%
MBGDM: $10^{-1}$
R: 99.74%
MBGDM: $10^{-2}$
UV
$1.337\times 10^{-3}$
MBGDM: $10^{-2}$
$5.259\times 10^{-6}$
MBGDM: $10^{-1}$
M: 99.69%
MBGDM: $10^{-2}$
E: 99.59%
MBGDM: $10^{-1}$
R: 99.56%
MBGDM: $10^{-1}$
UV-ReLU
$1.244\times 10^{-3}$
MBGDM: $10^{-1}$
$5.264\times 10^{-6}$
MBGDM: $10^{-1}$
M: 99.63%
MBGDM: $10^{-2}$
E: 99.68%
MBGDM: $10^{-1}$
R: 99.66%
MBGDM: $10^{-1}$
Figure SM10: Singular value spectrum corresponding to Table SM1.
B.3 Additional Details on the Pruning Experiment
The SVD-based pruning experiments yield conclusions consistent with the singular value decay pattern. UDV typically exhibits the fastest decay, leading to more compact models while maintaining performance, as shown in Figure SM13.
Figure SM11: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the regression tasks.
Figure SM12: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the MNIST dataset.
Figure SM13: The performance of SVD-based pruning. The index in x-axis represents the number of neurons in the diagonal layer after pruning. For the HPART and NYCTTD datasets, the validation loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$). Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale. The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice. For the MNIST dataset, the results show the validation accuracy after pruning the model.
We observed that the learning rate can have some impact on the singular value decay pattern. In particular, a very large learning rate may cause oscillations in both training and validation loss, yet may result in a rapid decay of the spectrum. Conversely, a small learning rate may lead to a less pronounced spectral decay but still can yield a comparable validation loss to that of the optimal learning rates. For example, the difference in validation losses between the Adam optimizer with learning rates of 1e-3 and 1e-4 was negligible, yet the singular value spectra differed (see Figure SM14). This discrepancy also affects the performance of the pruning experiments (see Figure SM15).
Figure SM14: Differences in singular value patterns across varying learning rates.
Figure SM15: Different learning rates lead different pruning performance. the validation loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$). Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale. The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice.
When the learning rate is close to the optimal value, a ‘stair-step’ pattern (see Figure SM16) can be observed on the loss or accuracy curve in the classification task. It could be attributed to the model continuously searching for a low-rank solution, increasing the rank gradually.
Figure SM16: The ‘stair-step’ accuracy curve in classification (RegNetX-32GF with Adam). The sub-figures in the second row correspond to the respective singular value spectra, indicating that UDV consistently seeks low-rank solutions regardless of the learning rate.
B.4 Incorporating ReLU into UDV
To explore whether non-linear activation functions, particularly ReLU, can be used in our UDV framework, we conducted an additional experiment.
Building on the standard ReLU, we first introduced norm constraints, $\sum_{j=1}^{m}\|\mathbf{u}_{j}\|_{2}^{2}\leq 1$ and $\sum_{j=1}^{m}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1$, to the layers both before and after ReLU, denoted as ReLU(constrained). Next, we integrated ReLU into the diagonal layer of UDV. When the non-negativity constraints are applied on the diagonal layer, we refer to the model as UDV-ReLU; otherwise, it is called UDV-ReLU-s.
We conducted a preliminary experiment to verify the feasibility of incorporating ReLU into the UDV, leaving a comprehensive study to future work. We derived the optimizers and learning rates from Tables SM2, SM3, SM4 and SM5, but only used the HPART dataset for regression and the MNIST for classification.
Figure SM17 and Figure SM18 showed that UDV-ReLU and UDV-ReLU-s exhibit a similar (or even more pronounced) decaying pattern in singular values as observed in the proposed UDV. The ‘stair-step’ accuracy curve was also observed in these models.
Figure SM17: ReLU in UDV: Singular value pattern on HPART dataset.
Figure SM18: ReLU in UDV: Validation accuracy (the first row) and the singular value pattern (the second row) on MNIST dataset. Faster convergence can be achieved by fine-tuning the learning rate.
B.5 Comparison of UDV and Three-Layer Fully Connected Networks
We extend the experiments to compare the UDV framework with three-layer fully connected neural networks. To highlight the critical role of the UDV structure and its constraints, we include the following models in the comparison: the proposed UDV model, a model based on the UDV structure but without the constraints (UDV unconstrained), a fully connected three-layer neural network (UFV), and the standard UV model as the baseline.
The results, also included in Tables SM2, SM3, SM4 and SM5, are shown in Figure SM19 and Figure SM20. These findings demonstrate that the singular value decay in UDV unconstrained, UFV, and UV is significantly slower than in the proposed UDV framework. This underscores the critical role of the diagonal layer and constraints in facilitating low-rank solution identification within the proposed structure.
Figure SM19: Comparison of singular value spectrum among UDV, UDV (unconstrained),UFV and UV on the regression tasks.
Figure SM20: Comparison of singular value spectrum among UDV, UDV (unconstrained), UFV and UV on the classification datasets (MNIST).
B.6 Comparing the Effects of UDV and Weight Decay Regularization
We conducted a preliminary experiment to compare the singular value decay in two-layer (UV) and three-layer (UFV) fully connected network blocks with weight decay regularization to that observed in the UDV framework. We focused on the regression task with the HPART dataset. For the regularization parameter ($\gamma$), we tested values $10^{-6}$, $10^{-5}$, $\dots$, $10^{-1}$.
Increasing $\gamma$ led to a faster decay in the singular value spectrum but at the cost of higher training and validation losses. Figure SM21 illustrates an example where we selected the largest $\gamma$ values for which the validation loss remained within 10% of the UDV baseline. The resulting singular value decay in the UV and UFV models was less pronounced than in the UDV model. Further increasing $\gamma$ to match or exceed the spectral decay of the UDV model resulted in significantly worse validation loss compared to the UDV network. The complete results are presented in Figures SM22 and SM23.
Figure SM21: Comparison of singular value spectrum among the proposed UDV, UFV and UV (with weight decay) on the HPART dataset. In addition to the UDV baseline, the UFV and UV models include the regularization parameter $\gamma$ and the relative change in validation performance. The relative change is defined as ($\frac{loss_{model}-loss_{baseline}}{loss_{baseline}}\times 100\%$), representing how much worse the model performs compared to the UDV baseline.
Figure SM22: Extension of Figure SM21. Comparison of singular value spectrum between the proposed UDV and UFV (with weight decay) on the HPART dataset. Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
Figure SM23: Extension of Figure SM21. Comparison of singular value spectrum between the proposed UDV and UV (with weight decay) on the HPART dataset. Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
Table SM2: Experiments using Adam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset
(Transferred model)
HPART
(×$\times 10^{-3}$)
NYCTTD
(×$\times 10^{-6}$)
MNIST
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)
LR: $10^{-6}$
UDV: –
UDV-s: –
UDV-v1: 99.56
UDV-v2: 99.53
UDV-ReLU: –
UDV(unconstrained): 99.53
UFV: 99.53
UV-ReLU: 99.59
UV: 99.54
M: 99.42
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: 99.25
UV-ReLU: 99.29
UV: 99.24
E: 98.67
UDV: –
UDV-s: –
UDV-v1: 99.46
UDV-v2: 99.48
UDV-ReLU: –
UDV(unconstrained): 99.48
UFV: 99.38
UV-ReLU: 99.41
UV: 99.36
R: 99.32
LR: $10^{-5}$
UDV: –
UDV-s: –
UDV-v1: 99.61
UDV-v2: 99.58
UDV-ReLU: –
UDV(unconstrained): 99.62
UFV: 99.60
UV-ReLU: 99.57
UV: 99.61
M: 99.59
UDV: –
UDV-s: –
UDV-v1: 99.49
UDV-v2: 99.53
UDV-ReLU: –
UDV(unconstrained): 99.54
UFV: 99.43
UV-ReLU: 99.52
UV: 99.52
E: 99.51
UDV: –
UDV-s: –
UDV-v1: 99.58
UDV-v2: 99.50
UDV-ReLU: –
UDV(unconstrained): 99.50
UFV: 99.58
UV-ReLU: 99.59
UV: 99.57
R: 99.59
LR: $10^{-4}$
UDV: 2.304
UDV-s: 1.912
UDV-v1: 1.823
UDV-v2: 1.738
UDV-ReLU: 2.731
UDV(unconstrained): 1.738
UFV: 1.351
UV-ReLU: 1.376
UV: 1.475
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.248
UDV-v2: 5.250
UDV-ReLU: 5.251
UDV(unconstrained): 5.250
UFV: 5.254
UV-ReLU: 5.263
UV: 5.275
UDV: 99.66
UDV-s: 99.58
UDV-v1: 99.65
UDV-v2: 99.69
UDV-ReLU: 99.63
UDV(unconstrained): 99.66
UFV: 99.69
UV-ReLU: 99.63
UV: 99.64
M: 99.65
UDV: –
UDV-s: –
UDV-v1: 99.63
UDV-v2: 99.58
UDV-ReLU: –
UDV(unconstrained): 99.63
UFV: 99.59
UV-ReLU: 99.54
UV: 99.56
E: 99.60
UDV: 99.55
UDV-s: 99.55
UDV-v1: 99.56
UDV-v2: 99.68
UDV-ReLU: 99.55
UDV(unconstrained): 99.59
UFV: 99.67
UV-ReLU: 99.64
UV:99.60
R: 99.59
LR: $10^{-3}$
UDV: 1.304
UDV-s: 1.316
UDV-v1: 1.267
UDV-v2: 1.268
UDV-ReLU: 1.320
UDV(unconstrained): 1.268
UFV: 1.318
UV-ReLU: 1.167
UV: 1.333
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.248
UDV-v2: 5.248
UDV-ReLU: 5.251
UDV(unconstrained): 5.248
UFV: 5.248
UV-ReLU: 5.306
UV: 5.251
UDV: 99.57
UDV-s: 99.59
UDV-v1: 99.57
UDV-v2: 99.58
UDV-ReLU: 99.51
UDV(unconstrained): 99.52
UFV: 99.59
UV-ReLU: 99.60
UV: 99.60
M: 99.63
UDV: 99.55
UDV-s: 99.59
UDV-v1: 99.59
UDV-v2: 99.57
UDV-ReLU: 99.54
UDV(unconstrained): 99.57
UFV: 99.54
UV-ReLU: 99.49
UV: 99.60
E: 99.61
UDV: 99.50
UDV-s: 99.49
UDV-v1: 99.47
UDV-v2: 99.47
UDV-ReLU: 99.42
UDV(unconstrained): 99.58
UFV: 99.41
UV-ReLU: 99.66
UV:99.46
R: 99.49
LR: $10^{-2}$
UDV: 1.877
UDV-s: 1.998
UDV-v1: 1.409
UDV-v2: 1.500
UDV-ReLU: 1.699
UDV(unconstrained): 1.402
UFV: 1.483
UV-ReLU: 1.467
UV: 1.430
UDV: 5.257
UDV-s:5.258
UDV-v1: 5.256
UDV-v2: 5.257
UDV-ReLU: 5.323
UDV(unconstrained): 5.263
UFV: 7.048
UV-ReLU: 5.323
UV: 7.369
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: 99.34
UDV-s: 99.56
UDV-v1: 99.48
UDV-v2: 99.44
UDV-ReLU: 96.95
UDV(unconstrained): 99.52
UFV: –
UV-ReLU: 99.37
UV: 99.32
E: 98.42
UDV: 99.53
UDV-s: 99.43
UDV-v1: 99.35
UDV-v2: 99.28
UDV-ReLU: 99.37
UDV(unconstrained): 99.10
UFV: –
UV-ReLU: 99.32
UV: 98.90
R: 99.39
LR: $10^{-1}$
UDV: 4.188
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: 18.26
UDV(unconstrained): –
UFV: 1.745
UV-ReLU: 42.01
UV: 1.614
UDV: 5.321
UDV-s: 114.6
UDV-v1: 23.12
UDV-v2: 21.22
UDV-ReLU: 5.323
UDV(unconstrained): 126.1
UFV: –
UV-ReLU: 5.323
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: 99.05
UDV-s: 95.62
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: 97.69
E: 99.10
UDV: 97.54
UDV-s: 99.31
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
R: 99.22
LR: $10^{0}$
UDV: 38.71
UDV-s: –
UDV-v1: 2.413
UDV-v2: 4.633
UDV-ReLU: 48.62
UDV(unconstrained): 14.05
UFV: –
UV-ReLU: 48.24
UV: –
UDV: 5.327
UDV-s: –
UDV-v1: 16.19
UDV-v2: –
UDV-ReLU: 5.323
UDV(unconstrained): –
UFV: –
UV-ReLU: 5.323
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
R: 95.65
LR: $2\times 10^{0}$
UDV: 60.46
UDV-s: –
UDV-v1: 6.651
UDV-v2: 7.366
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: 49.43
UV: –
UDV: 5.322
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: 5.323
UDV(unconstrained): –
UFV: –
UV-ReLU: 5.323
UV: –
LR: $3\times 10^{0}$
UDV: 106.9
UDV-s: –
UDV-v1: 6.606
UDV-v2: 7.398
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: 48.50
UV: –
UDV: 5.335
UDV-s: –
UDV-v1: 6.380
UDV-v2: –
UDV-ReLU: 5.323
UDV(unconstrained): –
UFV: –
UV-ReLU: 5.323
UV: –
LR: $5\times 10^{0}$
Table SM3: Experiments using NAdam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset
(Transferred model)
HPART
(×$\times 10^{-3}$)
NYCTTD
(×$\times 10^{-6}$)
MNIST
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)
LR: $10^{-6}$
UDV: –
UDV-s: –
UDV-v1: 99.56
UDV-v2: 99.53
UDV-ReLU: –
UDV(unconstrained): 99.53
UFV: 99.54
UV-ReLU: 99.59
UV: 99.53
M: 99.41
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: 99.26
UV-ReLU: 99.30
UV: 99.24
E: 98.67
UDV: –
UDV-s: –
UDV-v1: 99.49
UDV-v2: 99.46
UDV-ReLU: –
UDV(unconstrained): 99.46
UFV: 99.32
UV-ReLU: 99.40
UV: 99.32
R: 99.37
LR: $10^{-5}$
UDV: –
UDV-s: –
UDV-v1: 99.63
UDV-v2: 99.59
UDV-ReLU: –
UDV(unconstrained): 99.60
UFV: 99.62
UV-ReLU: 99.61
UV: 99.63
M: 99.62
UDV: –
UDV-s: –
UDV-v1: 99.50
UDV-v2: 99.52
UDV-ReLU: –
UDV(unconstrained): 99.52
UFV: 99.43
UV-ReLU: 99.51
UV: 99.52
E: 99.50
UDV: –
UDV-s: –
UDV-v1: 99.66
UDV-v2: 99.56
UDV-ReLU: –
UDV(unconstrained): 99.56
UFV: 99.54
UV-ReLU: 99.59
UV: 99.61
R: 99.67
LR: $10^{-4}$
UDV: 2.312
UDV-s: 1.916
UDV-v1: 1.837
UDV-v2: 1.752
UDV-ReLU: 2.665
UDV(unconstrained): 1.752
UFV: 1.381
UV-ReLU: 1.398
UV: 1.512
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.248
UDV-v2: 5.249
UDV-ReLU: 5.251
UDV(unconstrained): 5.249
UFV: 5.256
UV-ReLU: 5.258
UV: 5.275
UDV: 99.67
UDV-s: 99.62
UDV-v1: 99.67
UDV-v2: 99.61
UDV-ReLU: 99.63
UDV(unconstrained): 99.62
UFV: 99.68
UV-ReLU: 99.68
UV: 99.65
M: 99.59
UDV: –
UDV-s: –
UDV-v1: 99.60
UDV-v2: 99.66
UDV-ReLU: –
UDV(unconstrained): 99.60
UFV: 99.54
UV-ReLU: 99.53
UV: 99.53
E: 99.65
UDV: 99.52
UDV-s: 99.55
UDV-v1: 99.69
UDV-v2: 99.72
UDV-ReLU: 99.46
UDV(unconstrained): 99.63
UFV: 99.55
UV-ReLU: 99.63
UV: 99.64
R: 99.64
LR: $10^{-3}$
UDV: 1.638
UDV-s: 1.691
UDV-v1: 1.418
UDV-v2: 1.440
UDV-ReLU: 1.504
UDV(unconstrained): 1.437
UFV: 1.607
UV-ReLU: 1.367
UV: 1.654
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.248
UDV-v2: 5.248
UDV-ReLU: 5.251
UDV(unconstrained): 5.248
UFV: 5.248
UV-ReLU: 5.249
UV: 5.254
UDV: 99.53
UDV-s: 99.56
UDV-v1: 99.54
UDV-v2: 99.61
UDV-ReLU: 99.55
UDV(unconstrained): 99.66
UFV: 99.57
UV-ReLU: 99.63
UV: 99.59
M: 99.58
UDV: 97.56
UDV-s: 99.61
UDV-v1: 99.60
UDV-v2: 99.55
UDV-ReLU: 99.50
UDV(unconstrained): 99.55
UFV: 99.28
UV-ReLU: 99.55
UV: 99.55
E: 99.58
UDV: 99.45
UDV-s: 99.45
UDV-v1: 99.43
UDV-v2: 99.46
UDV-ReLU: 99.47
UDV(unconstrained): 99.50
UFV: 99.27
UV-ReLU: 99.53
UV:99.42
R: 99.65
LR: $10^{-2}$
UDV: 3.396
UDV-s: 3.287
UDV-v1: 2.297
UDV-v2: 2.315
UDV-ReLU: 2.403
UDV(unconstrained): 1.884
UFV: –
UV-ReLU: 1.954
UV: –
UDV: 5.258
UDV-s: 6.918
UDV-v1: 5.276
UDV-v2: 5.253
UDV-ReLU: 5.323
UDV(unconstrained): 5.253
UFV: 25.24
UV-ReLU: 5.323
UV: 6.262
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: 99.35
UDV-s: –
UDV-v1: 99.31
UDV-v2: 99.01
UDV-ReLU: 99.42
UDV(unconstrained): 98.78
UFV: –
UV-ReLU: 98.87
UV: 98.60
E: 99.26
UDV: 97.12
UDV-s: 99.28
UDV-v1: 99.37
UDV-v2: 99.38
UDV-ReLU: 99.49
UDV(unconstrained): 99.38
UFV: –
UV-ReLU: 99.28
UV: 98.91
R: 99.59
LR: $10^{-1}$
UDV: 13.69
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: 4.918
UV-ReLU: 41.61
UV: 1.863
UDV: 5.323
UDV-s: –
UDV-v1: 6.627
UDV-v2: –
UDV-ReLU: 5.323
UDV(unconstrained): –
UFV: –
UV-ReLU: 5.323
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: 97.54
E: 98.75
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
R: 99.19
LR: $10^{0}$
UDV: –
UDV-s: –
UDV-v1: 5.649
UDV-v2: 2.317
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: 45.77
UV: –
UDV: 5.328
UDV-s: 53.70
UDV-v1: 19.52
UDV-v2: –
UDV-ReLU: 5.323
UDV(unconstrained): –
UFV: –
UV-ReLU: 5.323
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
R: –
LR: $2\times 10^{0}$
UDV: –
UDV-s: –
UDV-v1: 3.075
UDV-v2: 2.828
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: 59.82
UV: –
UDV:5.377
UDV-s: –
UDV-v1: 6.011
UDV-v2: –
UDV-ReLU: 5.323
UDV(unconstrained): –
UFV: –
UV-ReLU: 5.323
UV: –
LR: $3\times 10^{0}$
UDV: –
UDV-s: –
UDV-v1: 3.882
UDV-v2: 3.128
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
UDV: 5.401
UDV-s: –
UDV-v1: 15.27
UDV-v2: –
UDV-ReLU: 5.323
UDV(unconstrained): –
UFV: –
UV-ReLU: 5.323
UV: –
LR: $5\times 10^{0}$
Table SM4: Experiments using MBGD optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset
(Transferred model)
HPART
(×$\times 10^{-3}$)
NYCTTD
(×$\times 10^{-6}$)
MNIST
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)
LR: $10^{-6}$
LR: $10^{-5}$
LR: $10^{-4}$
UDV: 46.90
UDV-s: 45.43
UDV-v1: 46.01
UDV-v2: 44.35
UDV-ReLU: 47.88
UDV(unconstrained): 44.35
UFV: 11.77
UV-ReLU: 12.59
UV: 11.98
UDV: 26.60
UDV-s: 40.06
UDV-v1: 49.35
UDV-v2: 83.91
UDV-ReLU: 20.56
UDV(unconstrained): 83.91
UFV: 86.68
UV-ReLU: –
UV: 71.72
LR: $10^{-3}$
UDV: 30.38
UDV-s: 21.39
UDV-v1: 23.65
UDV-v2: 15.77
UDV-ReLU: 39.89
UDV(unconstrained): 15.77
UFV: 6.719
UV-ReLU: 6.478
UV: 6.345
UDV: 9.407
UDV-s: 10.29
UDV-v1: 10.83
UDV-v2: 12.56
UDV-ReLU: 7.231
UDV(unconstrained): 12.56
UFV: 8.236
UV-ReLU: 12.67
UV: 6.763
UDV: –
UDV-s: –
UDV-v1: 98.44
UDV-v2: 98.32
UDV-ReLU: –
UDV(unconstrained): 98.27
UFV: 99.34
UV-ReLU: 99.24
UV: 99.19
M: 99.29
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: 98.15
UV-ReLU: 98.48
UV: 98.55
E: 98.74
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: 99.20
UV-ReLU: 99.16
UV: 99.17
R: 99.16
LR: $10^{-2}$
UDV: 6.030
UDV-s: 6.014
UDV-v1: 6.125
UDV-v2: 5.877
UDV-ReLU: 6.234
UDV(unconstrained): 5.877
UFV: 5.575
UV-ReLU: 2.253
UV: 2.251
UDV: 5.465
UDV-s: 5.421
UDV-v1: 5.514
UDV-v2: 5.532
UDV-ReLU: 5.431
UDV(unconstrained): 5.532
UFV: 5.293
UV-ReLU: 5.521
UV: 5.296
UDV: 99.38
UDV-s: 99.32
UDV-v1: 99.47
UDV-v2: 99.50
UDV-ReLU: 99.53
UDV(unconstrained): 99.58
UFV: 99.53
UV-ReLU: 99.53
UV: 99.50
M: 99.59
UDV: –
UDV-s: –
UDV-v1: 98.62
UDV-v2: 98.97
UDV-ReLU: –
UDV(unconstrained): 98.40
UFV: 99.05
UV-ReLU: 99.31
UV: 99.24
E: 99.43
UDV: 99.40
UDV-s: 99.46
UDV-v1: 99.37
UDV-v2: 99.34
UDV-ReLU: 99.47
UDV(unconstrained): 99.34
UFV: 99.36
UV-ReLU: 99.31
UV:99.38
R: 99.44
LR: $10^{-1}$
UDV: 1.398
UDV-s: 1.407
UDV-v1: 1.556
UDV-v2: 1.565
UDV-ReLU: 1.379
UDV(unconstrained): 1.565
UFV: 1.548
UV-ReLU: 1.493
UV: 1.583
UDV: 5.253
UDV-s: 5.250
UDV-v1: 5.256
UDV-v2: 5.252
UDV-ReLU: 5.277
UDV(unconstrained): 5.252
UFV: 5.255
UV-ReLU: 5.323
UV: 5.291
UDV: 99.59
UDV-s: 99.38
UDV-v1: 99.61
UDV-v2: 99.60
UDV-ReLU: 99.65
UDV(unconstrained): 99.59
UFV: 99.61
UV-ReLU: 99.60
UV: 99.63
M: 99.62
UDV: –
UDV-s: 99.36
UDV-v1: 99.43
UDV-v2: 99.53
UDV-ReLU: 95.21
UDV(unconstrained): 95.29
UFV: 99.55
UV-ReLU: 99.52
UV: 99.55
E: 99.67
UDV: 99.51
UDV-s: 99.57
UDV-v1: 99.60
UDV-v2: 99.62
UDV-ReLU: –
UDV(unconstrained): 99.61
UFV: 99.59
UV-ReLU: 99.59
UV: 98.56
R: 99.63
LR: $10^{0}$
UDV: 3.076
UDV-s: 2.870
UDV-v1: 1.935
UDV-v2: 1.830
UDV-ReLU: 4.573
UDV(unconstrained): 1.810
UFV: –
UV-ReLU: 48.62
UV: –
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.249
UDV-v2: 5.248
UDV-ReLU: 5.261
UDV(unconstrained): 5.248
UFV: 5.256
UV-ReLU: 5.271
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: 99.65
UDV-v2: 99.59
UDV-ReLU: –
UDV(unconstrained): 99.62
UFV: –
UV-ReLU: 99.64
UV: –
E: 99.55
UDV: –
UDV-s: –
UDV-v1: 99.67
UDV-v2: 99.69
UDV-ReLU: –
UDV(unconstrained): 99.57
UFV: –
UV-ReLU: 99.73
UV: 99.66
R: 99.66
LR: $2\times 10^{0}$
UDV: 6.244
UDV-s: 48.63
UDV-v1: 2.559
UDV-v2: –
UDV-ReLU: 47.50
UDV(unconstrained): –
UFV: –
UV-ReLU: 48.63
UV: –
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.249
UDV-v2: 5.249
UDV-ReLU: 5.259
UDV(unconstrained): 5.249
UFV: –
UV-ReLU: 5.293
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: 98.37
UDV-v2: 99.66
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: 98.88
UDV: –
UDV-s: –
UDV-v1: 99.66
UDV-v2: 99.55
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: 99.58
UV: 99.57
R: 99.56
LR: $3\times 10^{0}$
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: 48.62
UV: –
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.249
UDV-v2: 5.249
UDV-ReLU: 5.257
UDV(unconstrained): 5.250
UFV: –
UV-ReLU: 5.296
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: 99.07
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: 99.26
UDV: –
UDV-s: –
UDV-v1: 99.48
UDV-v2: 99.65
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: 99.57
UV: –
R: 99.57
LR: $5\times 10^{0}$
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: 99.33
UDV: –
UDV-s: –
UDV-v1: 99.55
UDV-v2: 99.41
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: 99.42
UV: –
R: 99.55
Table SM5: Experiments using MBGDM optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset
(Transferred model)
HPART
(×$\times 10^{-3}$)
NYCTTD
(×$\times 10^{-6}$)
MNIST
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)
LR: $10^{-6}$
LR: $10^{-5}$
LR: $10^{-4}$
UDV: 30.52
UDV-s: 21.50
UDV-v1: 23.79
UDV-v2: 15.84
UDV-ReLU: 40.00
UDV(unconstrained): 15.84
UFV: 6.729
UV-ReLU: 6.488
UV: 6.348
UDV: 9.413
UDV-s: 10.29
UDV-v1: 10.85
UDV-v2: 12.59
UDV-ReLU: 7.224
UDV(unconstrained): 12.59
UFV: 8.288
UV-ReLU: 12.68
UV: 6.776
LR: $10^{-3}$
UDV: 6.105
UDV-s: 6.080
UDV-v1: 6.178
UDV-v2: 5.929
UDV-ReLU: 6.341
UDV(unconstrained): 5.929
UFV: 2.590
UV-ReLU: 2.529
UV: 2.252
UDV: 5.457
UDV-s: 5.412
UDV-v1: 5.516
UDV-v2: 5.534
UDV-ReLU: 5.431
UDV(unconstrained): 5.534
UFV: 5.294
UV-ReLU: 5.520
UV: 5.296
UDV: 99.26
UDV-s: 99.38
UDV-v1: 99.39
UDV-v2: 99.52
UDV-ReLU: –
UDV(unconstrained): 99.51
UFV: 99.56
UV-ReLU: 99.46
UV: 99.54
M: 99.58
UDV: –
UDV-s: –
UDV-v1: 99.00
UDV-v2: 99.05
UDV-ReLU: –
UDV(unconstrained): 99.08
UFV: 99.18
UV-ReLU: 99.37
UV: 99.36
E: 99.45
UDV: 99.50
UDV-s: 99.43
UDV-v1: 99.39
UDV-v2: 99.37
UDV-ReLU: –
UDV(unconstrained): 99.40
UFV: 99.38
UV-ReLU: 99.36
UV: 99.30
R: 99.41
LR: $10^{-2}$
UDV: 1.357
UDV-s: 1.388
UDV-v1: 1.554
UDV-v2: 1.569
UDV-ReLU: 1.312
UDV(unconstrained): 1.569
UFV: 1.357
UV-ReLU: 1.314
UV: 1.337
UDV: 5.253
UDV-s: 5.250
UDV-v1: 5.256
UDV-v2: 5.252
UDV-ReLU: 5.276
UDV(unconstrained): 5.252
UFV: 5.254
UV-ReLU: 5.267
UV: 5.279
UDV: 99.67
UDV-s: 99.38
UDV-v1: 99.63
UDV-v2: 99.60
UDV-ReLU: 99.62
UDV(unconstrained): 99.59
UFV: 99.65
UV-ReLU: 99.63
UV: 99.69
M: 99.64
UDV: 99.48
UDV-s: 99.60
UDV-v1: 99.54
UDV-v2: 99.59
UDV-ReLU: 99.54
UDV(unconstrained): 99.55
UFV: 99.50
UV-ReLU: 99.53
UV: 99.50
E: 99.59
UDV: 99.74
UDV-s: 99.72
UDV-v1: 99.61
UDV-v2: 99.67
UDV-ReLU: 99.65
UDV(unconstrained): 99.63
UFV: 99.66
UV-ReLU: 99.61
UV:99.53
R: 99.60
LR: $10^{-1}$
UDV: 1.345
UDV-s: 1.339
UDV-v1: 1.313
UDV-v2: 1.302
UDV-ReLU: 1.318
UDV(unconstrained): 1.302
UFV: 1.358
UV-ReLU: 1.244
UV: –
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.249
UDV-v2: 5.248
UDV-ReLU: 5.259
UDV(unconstrained): 2.248
UFV: 5.251
UV-ReLU: 5.264
UV: 5.259
UDV: 99.61
UDV-s: 99.60
UDV-v1: 99.65
UDV-v2: 99.68
UDV-ReLU: –
UDV(unconstrained): 99.70
UFV: –
UV-ReLU: 99.63
UV: –
M: –
UDV: 99.63
UDV-s: 99.63
UDV-v1: 99.64
UDV-v2: 99.66
UDV-ReLU: 99.61
UDV(unconstrained): 99.59
UFV: 99.68
UV-ReLU: 99.68
UV: 99.59
E: 99.63
UDV: 99.60
UDV-s: 99.57
UDV-v1: 99.71
UDV-v2: 99.67
UDV-ReLU: –
UDV(unconstrained): 99.67
UFV: 99.63
UV-ReLU: 99.66
UV: 99.56
R: 99.67
LR: $10^{0}$
UDV: 123.6
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: 47.03
UDV(unconstrained): –
UFV: –
UV-ReLU: 48.62
UV: –
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.249
UDV-v2: 5.249
UDV-ReLU: 5.251
UDV(unconstrained): 5.249
UFV: –
UV-ReLU: 5.281
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: 99.54
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: 99.24
UDV: –
UDV-s: –
UDV-v1: 99.52
UDV-v2: 99.58
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: 99.53
UV: –
R: 99.34
LR: $2\times 10^{0}$
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: 48.62
UV: –
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.249
UDV-v2: 5.248
UDV-ReLU: 5.249
UDV(unconstrained): 5.248
UFV: –
UV-ReLU: 5.289
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: 98.95
UDV: –
UDV-s: –
UDV-v1: 99.31
UDV-v2: 99.37
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: 99.28
UV: –
R: 99.31
LR: $3\times 10^{0}$
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: 48.62
UDV(unconstrained): –
UFV: –
UV-ReLU: 48.62
UV: –
UDV: 5.248
UDV-s: 5.248
UDV-v1: 5.248
UDV-v2: 5.248
UDV-ReLU: 5.249
UDV(unconstrained): 5.248
UFV: –
UV-ReLU: 5.292
UV: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: 98.70
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
R: 99.12
LR: $5\times 10^{0}$
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
M: –
UDV: –
UDV-s: –
UDV-v1: –
UDV-v2: –
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: –
UV: –
E: –
UDV: –
UDV-s: –
UDV-v1: 99.55
UDV-v2: 99.41
UDV-ReLU: –
UDV(unconstrained): –
UFV: –
UV-ReLU: 99.42
UV: –
R: 99.55

### v2

@@
 Abstract
-Gradient descent for matrix factorization is known to exhibit an implicit bias toward approximately low-rank solutions.
+Gradient descent for matrix factorization exhibits an implicit bias toward approximately low-rank solutions.
 While existing theories often assume the boundedness of iterates, empirically the bias persists even with unbounded sequences.
-We thus hypothesize that implicit bias is driven by divergent dynamics markedly different from the convergent dynamics for data fitting.
-Using this perspective, we introduce a new factorization model: $X\approx UDV^{\top}$, where $U$ and $V$ are constrained within norm balls, while $D$ is a diagonal factor allowing the model to span the entire search space.
-Our experiments reveal that this model exhibits a strong implicit bias regardless of initialization and step size, yielding truly (rather than approximately) low-rank solutions.
-Furthermore, drawing parallels between matrix factorization and neural networks, we propose a novel neural network model featuring constrained layers and diagonal components.
-This model achieves strong performance across various regression and classification tasks while finding low-rank solutions, resulting in efficient and lightweight networks.
+This reflects a dynamic where factors develop low-rank structure while their magnitudes increase, tending to align with certain directions.
+To capture this behavior in a stable way, we introduce a new factorization model: $X\approx UDV^{\top}$, where $U$ and $V$ are constrained within norm balls, while $D$ is a diagonal factor allowing the model to span the entire search space.
+Experiments show that this model consistently exhibits a strong implicit bias, yielding truly (rather than approximately) low-rank solutions.
+Extending the idea to neural networks, we introduce a new model featuring constrained layers and diagonal components that achieves competitive performance on various regression and classification tasks while producing lightweight, low-rank representations.
 1 Introduction
 The Burer–Monteiro (BM) factorization (Burer & Monteiro, 2003) is a classical technique for obtaining low-rank solutions in optimization.
-It can be viewed as a simple neural network that uses a single layer of hidden neurons under linear activation.
-Indeed, say one has the factorization $X=UV^{T}$ where $U\in\mathbb{R}^{d\times r}$ and $V\in\mathbb{R}^{c\times r}$, then $U$ and $V$ can be thought of as the weights of the first and second layers, while $r$ represents the number of hidden neurons.
-However, despite the similarity suggested by this view, there is a clear distinction between BM factorization and neural networks about how the rank $r$ is chosen.
-In BM factorization, $r$ is typically chosen to be small, close to the rank of the desired solution.
-Neural networks, on the other hand, often succeed even in overparametrized settings where $r$ is large.
-Recent findings of implicit regularization in matrix factorization narrows the gap between these two perspectives.
-For instance, in an effort to explain the empirical success of overparametrized neural networks, Gunasekar et al.
+One can view it as a simple neural network that uses a single layer of hidden neurons under linear activation.
+Given the factorization $X=UV^{\top}$ where $U\in\mathbb{R}^{d\times r}$ and $V\in\mathbb{R}^{c\times r}$, $U$ and $V$ can be interpreted as the weights of the first and second layers, and $r$ as the number of hidden neurons.
+As a result, matrix factorization has become a convenient model for investigating the principles underlying the empirical success of neural networks.
+For instance, Gunasekar et al.
 (2017) demonstrate that gradient descent (with certain parameter selection) on BM factorization tends to converge toward approximately low-rank solutions even when $r=d$.
-Based on this observation, they conjecture that “with small enough step sizes and initialization close enough to the origin, gradient descent on full-dimensional factorization converges to the minimum nuclear norm solution.”
-In a follow-up work, however, Razin & Cohen (2020) presented a counter-example demonstrating that implicit regularization in BM factorization cannot be explained by minimal nuclear norm, or in fact any norm.
-Specifically, they showed that there are instances where the gradient method applied to BM factorization yields a diverging sequence, and all norms thus grow toward infinity.
+Based on this observation, they conjecture that “with small enough step-sizes and initialization close enough to the origin, gradient descent on full-dimensional factorization converges to the minimum nuclear norm solution.”
+In a follow-up work, Razin & Cohen (2020) present a counterexample demonstrating that implicit regularization in BM factorization cannot be explained by minimal nuclear norm, or in fact any norm.
+Specifically, they present instances where the gradient method applied to BM factorization yields a diverging sequence, and all norms thus grow toward infinity.
 Intriguingly, despite this divergence, they found that the rank of the estimate decreases toward its minimum.
-Although this phenomenon might seem surprising initially, it is not uncommon for diverging sequences to follow a structured path.
+In fact, it is not uncommon for diverging sequences to follow a structured path.
 A prime example is the Power Method, the fundamental algorithm for finding the largest eigenvalue and eigenvector pair of a matrix.
@@
 This process amplifies the component of the vector that aligns with the direction of the dominant eigenvector more than the other components, progressively leading $x_{k}$ to align with this eigenvector.
-In practical implementations, $x_{k}$ is scaled after each iteration to avoid numerical issues from divergence.
-This perspective serves as the foundational motivation guiding our approach.
-Specifically, our key insight is that the implicit regularization in BM factorization (and neural networks) is driven by divergent dynamical behavior.
-This is markedly different from the standard (convergent) optimization dynamics helping with the data fitting.
-In this context, we hypothesize that these forces do not merely coexist but actively compete, influencing model behavior and performance in fundamentally conflicting ways.
-Our main goal in the development of this paper is to devise an approach that unravels these competing forces.
+In practice, $x_{k}$ is scaled after each iteration to avoid numerical issues from divergence.
+This perspective underpins our approach.
+We interpret implicit bias in BM factorization as arising from growth dynamics that amplify certain directions while suppressing others, thereby inducing low-rank structure.
+To capture this effect in a stable way, we propose a new factorization model that disentangles structured bias from instability.
 1.1 Overview of main contributions
   • $\blacksquare$
-    A novel formulation for matrix factorization.
-We model $X=UDV^{\top}$, where $U$ and $V$ are constrained within Frobenius norm balls.
-Projection onto this ball results in a scaling step similar to the Power Method.
+    A novel matrix factorization formulation.
+We introduce a new factorization model, $X=UDV^{\top}$, where $U$ and $V$ are constrained within Frobenius norm balls.
+Projection onto this ball results in a scaling step.
 The middle term $D$ is a diagonal matrix that allows the model to explore the entire search space despite $U$ and $V$ being bounded.
-    Empirically we demonstrate that the gradient method applied to the proposed formulation exhibits a pronounced implicit bias toward low-rank solutions.
-We compare our formulation against standard BM factorization with two unconstrained factors.
-Specifically, we investigate key factors such as step size and initialization, which prior work suggests might be contributing to implicit bias.
+[^1] The reader may notice a “syntactic” similarity with singular value decomposition; except using simpler Frobenius norm constraints on $U$ and $V$ instead of orthogonality.
+    Through extensive empirics we demonstrate that the gradient method applied to the proposed formulation exhibits a pronounced implicit bias toward low-rank solutions.
+We compare our formulation against the standard BM factorization with two unconstrained factors.
+We investigate how step size and initialization influence implicit bias, building on prior work that points to these as potential contributors.
 We find that our factorization approach largely obviates the need to rely on these conditions: it consistently finds truly (rather than approximately) low-rank solutions across a wide range of initializations and step-sizes in our experiments.
-We believe these findings should be of broader interest to research on implicit bias.
   • $\blacksquare$
     A novel neural network architecture.
-Motivated by the strong bias for low-rank solutions of the proposed factorization, we subsequently extend it to deep neural networks.
-Specifically, we do so by adding constrained layers and diagonal components.
-We demonstrate numerically that this constrained model performs on par with, or even better than, the standard architecture across various regression and classification tasks.
-Importantly, even here, our approach exhibits bias towards low-rank solutions, resulting in a natural pruning procedure to obtain compact, lightweight networks without compromising performance.
-2 Related Work
+Motivated by the strong bias for low-rank solutions of the proposed factorization, we extend it to deep neural networks.
+We do so by adding constrained layers and diagonal components.
+We show that this new model achieves competitive performance with standard architectures across various regression and classification tasks.
+Importantly, our approach exhibits bias towards low-rank solutions, resulting in a natural pruning procedure that delivers compact, lightweight networks without compromising performance.
+1.2 Related Work
 Burer-Monteiro factorization.
-BM factorization is initially proposed for solving semi-definite programs (Burer & Monteiro, 2003, 2005; Boumal et al., 2016, 2020; Cifuentes, 2021) and is recognized for its efficiency in addressing low-rank matrix optimization problems, see (Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017, 2018; Hsieh et al., 2018; Sahin et al., 2019; Lee et al., 2022; Yalçın et al., 2023) and the references therein.
-Building on the connections between training problems for (non-linear) two-layer neural networks and convex (copositive) formulations (see (Pilanci & Ergen, 2020; Ergen & Pilanci, 2020; Sahiner et al., 2021) and the references therein), Sahiner et al.
-(2023) recently introduced BM factorization to solve convex formulations for various neural network architectures, including fully connected networks with ReLU and gated ReLU (Fiat et al., 2019) activations, two-layer convolutional neural networks (LeCun et al., 1989a), and self-attention mechanisms based on transformers (Vaswani et al., 2017).
+BM factorization was proposed for solving semidefinite programs (Burer & Monteiro, 2003, 2005) and has been recognized for its efficiency in addressing low-rank optimization problems (Boumal et al., 2016; Park et al., 2018).
+Building on the connections between matrix factorization and training problems for two-layer neural networks, BM has served as foundational model for understanding implicit bias and developing theoretical insights.
 Implicit regularization.
-While neural network-based systems are rapidly emerging as a dominant technology, the mechanisms underlying their generalization capabilities are yet to be fully understood.
-One promising line of research aims to explain this success through the concept of ‘implicit regularization,’ which is induced by the optimization methods and formulations used during neural network training (Neyshabur et al., 2014, 2017; Neyshabur, 2017).
-Several studies have explored matrix factorization to investigate implicit bias in linear neural networks (Gunasekar et al., 2017; Arora et al., 2018; Razin & Cohen, 2020; Belabbas, 2020; Li et al., 2021).
-Le & Jegelka (2022) extended these results to the final linear layers of nonlinear ReLU-activated feedforward networks with fully connected layers and skip connections.
-More recently Timor et al.
-(2023) investigated implicit regularization in ReLU networks.
+One promising line of research that aims to explain the successful generalization abilities of neural networks is that of ‘implicit regularization’ induced by the optimization methods and architectures (Neyshabur et al., 2014, 2017; Neyshabur, 2017).
+Several studies explore matrix factorization to investigate implicit bias (Gunasekar et al., 2017; Arora et al., 2018; Razin & Cohen, 2020; Belabbas, 2020; Li et al., 2021).
 Much of the existing work focuses on gradient flow dynamics in the limit of infinitesimal learning rates.
-In particular, Gidel et al.
-(2019) examined discrete gradient dynamics in two-layer linear neural networks, showing that the dynamics progressively learn solutions of reduced-rank regression with a gradually increasing rank.
+Exceptionally, Gidel et al.
+(2019) examine discrete gradient dynamics in two-layer linear neural networks, showing that the dynamics progressively learn solutions of reduced-rank regression with a gradually increasing rank.
 Constrained neural networks.
-Regularizers are frequently used in neural network training to prevent overfitting and improve generalization (Goodfellow et al., 2016), or to achieve structural benefits such as sparse and compact network architectures (Scardapane et al., 2017).
-However, it is conventional to apply these regularizers as penalty functions in the objective rather than as hard constraints, addressing them within gradient-based optimization via their (sub)gradients.
+Regularizers are frequently used in neural network training to prevent overfitting and improve generalization, or to achieve structural benefits such as sparse and compact network architectures (Scardapane et al., 2017).
+However, it is conventional to apply these regularizers as penalty functions in the objective rather than constraints.
 This approach is likely favored due to the ease of implementation, as pre-built functions are readily available in common neural network packages.
-Several independent studies have also applied proximal methods across different networks and applications, finding that proximal-based methods tend to yield solutions with more pronounced structures (Bai et al., 2019; Yang et al., 2020; Yun et al., 2021; Yurtsever et al., 2021; Yang et al., 2022).
-Structural regularization in the form of hard constraints, however, appears to be rare in neural network training.
-One notable exception is in the context of neural network training with the Frank-Wolfe algorithm (Pokutta et al., 2020; Zimmer et al., 2022; Macdonald et al., 2022), where constraints are necessary due to the algorithm’s requirement for a bounded domain.
-There are other use cases of constraints in neural networks.
-For instance, constraints can be applied to ensure adherence to real-world conditions in applications where they are necessary (Pathak et al., 2015; Márquez-Neila et al., 2017; Jia et al., 2017; Kervadec et al., 2019; Weber et al., 2021), or to incorporate physical laws in physics-informed neural networks (Raissi et al., 2019; Lu et al., 2021; Patel et al., 2022).
-In these cases, the feasible set is typically complex and difficult to project onto; therefore, optimization algorithms like primal-dual methods or augmented Lagrangian techniques are used.
-Constraints are also present in lifted neural networks, a framework where the training problem is reformulated in a higher-dimensional ‘lifted’ space.
-In this space, conventional activation functions like ReLU, used in the original formulation, are expressed as constraints (Askari et al., 2018; Sahiner et al., 2021; Bartan & Pilanci, 2021; Prakhya et al., 2024).
+Regularization in the form of constraints appears to be rare in neural network training.
+One notable exception is in the context of neural network training with the Frank-Wolfe algorithm (Pokutta et al., 2020; Zimmer et al., 2022; Macdonald et al., 2022).
+Recently, Pethick et al.
+(2025) revealed parallels between Frank-Wolfe on constrained networks and algorithms that post-process update steps, such as Muon (Jordan et al., 2024), which achieves state-of-the-art results on nanoGPT by orthogonalizing the update directions before applying them.
 Pruning.
-Neural networks are often trained in an over-parameterized regime to enhance generalization and avoid getting stuck in poor local minima.
-However, these models can suffer from excessive memory and computational resource demands, making them less efficient for deployment in real-world applications (Chang et al., 2021).
-Several compression techniques have been studied in the literature, including parameter quantization (Krishnamoorthi, 2018), knowledge distillation (Gou et al., 2021), and pruning.
+Neural networks are often overparameterized, which can enhance generalization and avoid poor local minima.
+But such models then suffer from excessive memory and computational demands, making them less efficient for deployment in real-world applications (Chang et al., 2021).
 Pruning reduces the number of parameters, resulting in more compact and efficient models that are easier to deploy.
-The literature on pruning is extensive, with diverse methods proposed with distinct characteristics.
-Various criteria are used to determine which weights to prune, including second-order derivative-based methods (LeCun et al., 1989b; Hassibi & Stork, 1992), magnitude-based pruning (Janowsky, 1989; Han et al., 2015), saliency heuristics (Mozer & Smolensky, 1988; Lee et al., 2018), and matrix or tensor factorization-based techniques (Xue et al., 2013; Sainath et al., 2013; Jaderberg et al., 2014; Lebedev et al., 2015; Swaminathan et al., 2020), among others.
-A comprehensive review of pruning methods is beyond the scope of this paper due to space limitations and the diversity of approaches.
-For more detailed reviews, we refer to (Reed, 1993; Blalock et al., 2020; Cheng et al., 2024), and the references therein.
+A comprehensive review on pruning is beyond the scope of this paper due to space limitations and the diversity of approaches.
+We refer to (Reed, 1993; Blalock et al., 2020; Cheng et al., 2024) and the references therein for detailed reviews.
+Pruning by singular value thresholding has recently shown promising results, particularly in natural language processing (Chen et al., 2021), and is often used along with various enhancements such as importance weights and data whitening for effective compression of large language models (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024).
 2 Matrix Factorization with a Diagonal Component
-Consider the matrix sensing problem, where the goal is to recover a positive semidefinite (PSD) matrix $X\in\mathbb{S}_{+}^{d\times d}$ from a set of linear measurements $\smash{b=\mathcal{A}(X)\in\mathbb{R}^{n}}$.
-We define $\smash{\mathcal{A}:\mathbb{R}^{d\times d}\rightarrow\mathbb{R}^{n}}$ through symmetric measurement matrices $\smash{A_{1},\ldots,A_{n}\in\mathbb{S}^{d\times d}}$, such that $\mathcal{A}(X)=[\langle A_{1},X\rangle~{}\cdots~{}\langle A_{n},X\rangle]^{\top}$ and $\mathcal{A}^{\top}y=\sum_{i=1}^{n}y_{i}A_{i}.$ We particularly focus on the data-scarce setting where $n\ll d^{2}$.
-A popular example is the matrix completion problem, where the goal is to reconstruct a matrix $X$ from a subset of its entries.
-This problem is inherently under-determined; however, successful recovery is possible if $X$ is known to be low-rank (Candes & Recht, 2012).
-Remark 1.
-We focus on the recovery problem of a PSD matrix for simplicity; note that a non-square matrix sensing problem can also be reformulated as a PSD matrix sensing problem through a simple transformation (Park et al., 2017).
-The problem described above can be cast as the following rank-constrained optimization problem:
-$\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad f(X):=\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{subj.\ to}\quad\mathrm{rank}(X)\leq r.$ (1)
-Although rank-constrained matrix optimization problems are typically NP-hard, various methods have been developed to provide practical approximations.
-Examples include hard thresholding algorithms (Jain et al., 2010; Goldfarb & Ma, 2011; Kyrillidis & Cevher, 2014), convex relaxation techniques (Candes & Recht, 2012; Recht et al., 2010), and BM factorization (Burer & Monteiro, 2003; Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017).
-The main idea behind BM factorization is to reparametrize the decision variable $X$ as $UU^{\top}$, where the factor $U\in\mathbb{R}^{d\times r}$, and $r$ is a positive integer that controls the rank of the resulting product.
-Problem (1) can then be reformulated as:
+Consider the problem of reconstructing a matrix $\smash{X\in\mathbb{R}^{d\times c}}$ from a set of linear measurements $\smash{b=\mathcal{A}(X)\in\mathbb{R}^{n}}$.
+We focus on recovering a PSD matrix, without loss of generality since the general case can be reformulated as a PSD matrix sensing problem (Park et al., 2017).
+We particularly focus on the data-scarce setting where $n\ll d^{2}$, a notable example is matrix completion.
+This problem is inherently under-determined, but successful recovery is possible if $X$ is low-rank (Candes & Recht, 2012).
+One popular approach is the BM factorization, which reparametrizes the decision variable $X$ as $UU^{\top}$ with $U\in\mathbb{R}^{d\times r}$, and $r$ is a positive integer that controls the rank of the resulting product.
+Then, the problem can be formulated as:
 $\min_{U\in\mathbb{R}^{d\times r}}\quad\frac{1}{2}\|\mathcal{A}(UU^{\top})-b\|_{2}^{2}.$ (1)
-Despite the fact that finding the global minimum of (2) remains challenging, a local solution can be approximated using the gradient method (Lee et al., 2016).
-Initializing at $U_{0}\in\mathbb{R}^{d\times r}$, perform the iteration:
-$U_{k+1}=U_{k}-\eta\nabla_{U}f(U_{k}U_{k}^{\top}),$ (3)
-where $\eta>0$ is the step-size, and the gradient $\nabla_{U}f$ is computed as
-$\nabla_{U}f(UU^{\top})=2\nabla f(UU^{\top})U=2\mathcal{A}^{\top}(\mathcal{A}(UU^{\top})-b)U.$
-Selecting the factorization rank $r$ is a critical decision that might impact the solution.
+While finding the global minimum is intractable, a local solution can be found by using gradient descent (Lee et al., 2016). Initializing at some $U_{0}\in\mathbb{R}^{d\times r}$, perform:
+$U_{k+1}=U_{k}-\eta\nabla_{U}f(U_{k}U_{k}^{\top}),$
+where $\eta>0$ is the step-size, and the gradient is computed as $\nabla_{U}f(UU^{\top})=2\nabla f(UU^{\top})U$.
+Selecting the factorization rank $r$ is a critical decision.
 A small $r$ may lead to spurious local minima, resulting in inaccurate outcomes (Waldspurger & Waters, 2020).
 Conversely, a large $r$ might weaken rank regularization, rendering the problem underdetermined.
-Conventional wisdom in BM factorization suggests finding a moderate choice that balances these two extremes.
-However, a key observation in (Gunasekar et al., 2017) is that the gradient method applied to (2) exhibits a tendency towards approximately low-rank solutions (i.e., a decaying singular value spectrum) even when $r=d$.
-Below, we restate their conjecture:
-Conjecture in (Gunasekar et al., 2017).
-Suppose gradient flow (i.e., gradient descent with an infinitesimally small step-size) is initialized at a full-rank matrix arbitrarily close to the origin.
-If the limit of the gradient flow, $X_{\text{GF}}=UU^{\top}$, exists and is a global optimum of (1) with $\mathcal{A}(X_{\text{GF}})=b$, then $X_{\text{GF}}$ is the minimal nuclear-norm solution to (1).
-Remark 2.
-This conjecture outlines three conditions for implicit bias: a small step-size, near-origin full-rank initialization, and consistent measurements ($\mathcal{A},b$).
-However, these conditions are often unconventional.
-For example, small step-sizes are typically avoided because they slow down convergence.
-Similarly, near-origin initialization is counterintuitive, as the origin is a trivial spurious stationary point of (2).
+Conventional wisdom suggests finding a moderate compromise between these two extremes.
+However, a key observation in (Gunasekar et al., 2017) is that the gradient method applied to (1) exhibits a tendency towards approximately low-rank solutions even when $r=d$.
 2.1 The Proposed Factorization
 We propose reparameterizing $X=UDU^{\top}$, where $U\in\mathbb{R}^{d\times r}$ is constrained to have a bounded norm, and $D\in\mathbb{R}^{r\times r}$ is a non-negative diagonal matrix:
-$\min_{\begin{subarray}{c}U\in\mathbb{R}^{d\times r}\\ D\in\mathbb{R}^{r\times r}\end{subarray}}~{}\frac{1}{2}\|\mathcal{A}(UDU^{\top})-b\|_{2}^{2}\quad\mathrm{s.t.}\quad\|U\|_{F}\leq\alpha,~{}~{}D_{ii}\geq 0,~{}~{}D_{ij}=0,~{}~{}\forall i~{}\text{and}~{}\forall j\neq i,$ (4)
+$\min_{\begin{subarray}{c}U\in\mathbb{R}^{d\times r}\\ D\in\mathbb{R}^{r\times r}\end{subarray}}~\frac{1}{2}\|\mathcal{A}(UDU^{\top})-b\|_{2}^{2}\quad\mathrm{s.t.}\quad\|U\|_{F}\leq\alpha,~~D_{ii}\geq 0,~~D_{ij}=0,\quad\forall i~\text{and}~\forall j\neq i$ (2)
 where $\alpha>0$ is a model parameter. When the problem is well-scaled, for instance through basic preprocessing with data normalization, we found that $\alpha=1$ is a reasonable choice.
-Placing in multiple factors and with constraints, we perform projected-gradient updates on $U$ and $D$ with step-size $\eta>0$:
+We perform simultaneous projected-gradient updates on $U$ and $D$ with step-size $\eta>0$:
 $U_{k+1}$ $=\Pi_{U}\left(U_{k}-\eta\nabla_{U}f(U_{k}D_{k}U_{k}^{\top})\right)$ (3)
@@
 where $\Pi_{U}$ and $\Pi_{D}$ are projections for the constraints in (2); while the gradients are
-$\nabla_{U}f(UDU^{\top})$ $=2\nabla f(UDU^{\top})UD,$
+$\nabla_{U}f(UDU^{\top})=2\nabla f(UDU^{\top})UD$
 $\nabla_{D}f(UDU^{\top})=U^{\top}\nabla f(UDU^{\top})U.$
-3.2 Numerical Experiments on Matrix Factorization
-We present numerical experiments comparing the empirical performance of the proposed approach with the classical BM factorization.
-Specifically, we examine the impact of initialization and step-size on the singular value spectrum of the resulting solution.
-We set up a synthetic matrix completion problem to recover a PSD matrix $\smash{X_{\natural}=U_{\natural}U_{\natural}^{\top}}$, where the entries of $\smash{U_{\natural}\in\mathbb{R}^{100\times 3}}$ are independently and identically (iid) drawn from the standard Gaussian distribution.
+2.2 Experiments on Matrix Factorization
+We present numerical experiments that compare the proposed approach with the classical BM factorization on synthetic matrix completion and on a real Fourier ptychography phase retrieval task.
+2.2.1 Matrix Completion
+We set up a synthetic matrix completion problem to recover a PSD matrix $\smash{X_{\natural}=U_{\natural}U_{\natural}^{\top}}$, where the entries of $\smash{U_{\natural}\in\mathbb{R}^{100\times 3}}$ are drawn independently from $N(0,1)$.
 We randomly sample $n=900$ entries of $X_{\natural}$ and store them in the vector $b\in\mathbb{R}^{n}$.
 The goal is to recover $X_{\natural}$ from $b$ by solving problems (1) and (2).
-For initialization, we generate $U_{0}\in\mathbb{R}^{d\times d}$ with entries drawn iid from standard Gaussian distribution and rescale it to have the Frobenius norm $\xi>0$ (we investigate the impact of $\xi$).
-We initiate $D_{0}$ from the identity matrix.
+Although the ground-truth rank is three, we adopt a full-rank factorization with $r=d$ in order to study the implicit bias.
+For initialization, we generate $U_{0}\in\mathbb{R}^{d\times d}$ with entries drawn independently from $N(0,1)$; and rescale it to have Frobenius norm $\xi>0$, which sets its initial distance from the origin.
+For the UDU formulation, we set $D_{0}=I$ and the constraint parameter $\alpha=1$.
+Our experiments focus on how initialization ($\xi$) and step size ($\eta$) affect the singular value spectrum of the solutions.
+We first investigate the impact of step-size.
+To this end, we fix $\xi=10^{-2}$ and test different values of $\eta$.
 The results are shown in Figure 1.
-First, we examine the impact of step-size.
-To this end, we fix $\xi=10^{-2}$ and test different values of $\eta$.
 In the left panel, we plot the objective residual as a function of iterations.
 As expected, we observe that a smaller step-size slows down convergence.
-In the right panel, we plot the singular value spectrum of the results attained after $10^{6}$ iterations.
-We observe no direct connection between step-size and implicit bias in BM factorization.
-Impact of step-size ($\eta$), in noiseless setting, with fixed initialization.
-Impact of initial distance to origin $(\xi)$, in noiseless setting, with fixed step-size.
+The right panel presents the singular value spectra after $10^{6}$ iterations, revealing no clear correlation between step size and the spectral decay of the solutions.
+Notably, while the classical BM factorization finds an approximately low-rank solution, the proposed factorization converges to a solution of true rank three, matching the ground truth.
+Impact of step-size ($\eta$), with fixed initialization.
+Impact of initial distance to origin $(\xi)$, with fixed step-size.
 Figure 1: Impact of step-size and initialization on implicit bias.
@@
 In all cases, UDU produces truly low-rank solutions, whereas the classical approach results in approximate low-rank structures.
-Next, we investigate the impact of initialization.
-We fix the step-size at $\eta=10^{-2}$ and evaluate the effect of varying $\xi$.
-We observe a correlation between the implicit bias of the BM factorization and $\xi$, which determines the initial distance from the origin.
-Initializing closer to the origin in the classical BM factorization yields solutions with a faster spectral decay.
-Notably, the UDU factorization demonstrates a strong implicit bias toward truly low-rank solutions, regardless of the choice of $\eta$ or $\xi$.
-We provide additional experiments on matrix factorization problems in the Appendices.
-Specifically, Section A.1 considers the matrix completion problem with noisy measurements, where $b$ is perturbed with Gaussian noise.
-The results remain consistent with the noiseless case: the UDU model exhibits an implicit bias toward truly low-rank solutions, while the classical BM factorization yields approximately low-rank solutions, with spectral decay strongly influenced by initialization.
-In Section A.2, we present the evolution of the factors $U$ and $D$ over the iterations.
-Interestingly, the UDU model produces rank-revealing solutions, with the columns of $U$ converging to zero in certain directions.
-Specifically, $U$ tends to grow along certain directions, while the rescaling induced by projection onto the bounded constraint shrinks other coordinates.
-This behavior aligns with the mechanism of the power method and offers insights into its connection with divergent forces.
-Additionally, we present numerical experiments on a matrix sensing problem arising in phase retrieval image recovery in Section A.3.
-As before, the UDU framework consistently promotes low-rank solutions, and this structural bias significantly enhances the quality of the recovered image, as demonstrated by our results.
-4 Feedforward Neural Networks with Diagonal Hidden Layers
+Next, we examine the impact of initialization.
+We fix the step-size at $\eta=10^{-2}$ and evaluate the effect of $\xi$.
+In the classical BM factorization, initializing closer to the origin leads to solutions with faster spectral decay, consistent with the observations of Gunasekar et al.
+(2017).
+In contrast, the proposed UDU factorization exhibits a pronounced implicit bias toward exactly low-rank solutions, independent of the choice of $\xi$.
+We also conducted the same experiments under Gaussian noise in Section A.1, and the results are qualitatively consistent.
+The UDU model continues to exhibit an implicit bias toward truly low-rank solutions, while the classical method gets approximately low-rank solutions.
+Exact rank-3 recovery is no longer possible in the noisy setting, but the UDU factorization still generates a truly low-rank solution, at a slightly higher rank.
+2.2.2 Fourier Ptychography
+We consider a phase retrieval problem that arises in Fourier ptychography (FP), an imaging technique that reconstructs the phase of a complex-valued image (the sample’s transmission function) from multiple Fourier measurements under illuminations from different angles.
+We use real measurements acquired from a functional FP system applied to a human blood cell sample, obtained from Horstmeyer et al.
+(2015).
+The target is a $160\times 160$ complex-valued image, vectorized into $d=160^{2}=25^{\prime}600$ dimensions and represented by a vector $x_{\natural}\in\mathbb{C}^{d}$.
+The measurements are modeled as $b_{i}\approx|\langle a_{i},x_{\natural}\rangle|^{2}$ for $i=1,\dots,n$, where the approximation accounts for experimental noise from sensor limitations, illumination variability, and other imperfections.
+The sensing vectors $a_{i}\in\mathbb{C}^{d}$ are derived from windowed Fourier transforms under different illumination patterns.
+The dataset consists of $29$ illumination patterns, each recorded as an $80\times 80$ intensity image, yielding a total of $n=185^{\prime}600$ measurements.
+FP is a nonconvex quadratic inverse problem, which admits a lifted formulation as a PSD matrix recovery problem with a rank-1 constraint.
+The lifted formulation corresponds to recovering a $25^{\prime}600\times 25^{\prime}600$ complex PSD matrix from the measurements, which makes solving at full-rank factorization (with $r=d$) computationally infeasible.
+Instead, we set the factorization rank to $r=20$ and compare the proposed UDU formulation with the standard BM factorization.
+Both methods are run for $10^{4}$ iterations with step size $0.1$, initialized from the same random $U_{0}$ with i.i.d.
+Gaussian complex entries normalized in Frobenius norm.
+For UDU, we set $D_{0}=I$ and $\alpha=1$.
+Once solved, we recover the image estimate by extracting the dominant singular vector from the ($d\times r$) dimensional factors.
+The results are shown in Figure 2.
+The UDU factorization successfully constructs the image, whereas BM produces a heavily corrupted reconstruction.
+An inspection of the singular value spectrum reveals that UDU effectively converges to a rank-1 solution, while BM yields a spectrum with slow decay, explaining the difference in reconstruction quality.
+Importantly, UDU achieves this from a random initialization, in contrast to existing nonconvex phase retrieval methods that typically require carefully designed spectral initialization procedures (Candes et al., 2015; Bian et al., 2015).
+Figure 2: [Left] Singular value spectra of the solutions obtained by BM and UDU after $10^{4}$ iterations with step size $0.1$.
+[Middle] The corresponding reconstructed image from the BM factorization exhibits artifacts.
+[Right] The UDU factorization produces a clean recovery.
+Additional details are provided in Section A.2.
+There, we show reconstructions at intermediate iterations, and experiments on smaller synthetic phase retrieval problems using full-rank factorizations ($r=d$).
+In all cases, the UDU framework promotes low-rank solutions, and this structural bias improves phase recovery.
+2.3 Theoretical Insights into the Inner Workings and Implicit Bias
+A fixed-point analysis of the proposed method provides valuable insights into its inner workings.
+Lemma 1 (Fixed-point characterization).
+Define the update variables before projection as $\bar{U}=U-2\eta\nabla f(X)UD$ and $\bar{D}=D-\eta U^{\top}\nabla f(X)U$, with $X=UDU^{\top}\!$.
+Suppose $(U,D)$ is a fixed point of the algorithm in (3), and let $u_{j}$ denote the $j^{\text{th}}$ column of $U$ and $\lambda_{j}$ the $j^{\text{th}}$ diagonal entry of $D$.
+Then,
+  1. (a)
+    If $\|\bar{U}\|_{F}\leq\alpha$, then $\nabla f(X)u_{j}\lambda_{j}=0$ for all $j$.
+  2. (b)
+    If $\|\bar{U}\|_{F}>\alpha$, then there exists some $\beta>0$ such that $\nabla f(X)u_{j}\lambda_{j}=-\beta u_{j}$ for all $j$.
+  3. (c)
+    If $\lambda_{j}=0$, then $\smash{u_{j}^{\top}\nabla f(X)u_{j}\geq 0}$.
+  4. (d)
+    If $\lambda_{j}>0$, then $\smash{u_{j}^{\top}\nabla f(X)u_{j}=0}$.
+The proof is given in Section A.4.
+At first glance, condition $(b)$ might suggest that choosing a smaller $\alpha$ leads to fixed points where the columns of $U$ align with the negative eigenvectors of $\nabla f(X)$.
+However, the following proposition shows that no valid fixed point can satisfy $\|\bar{U}\|_{F}>\alpha$.
+Proposition 1 (Exclusion of spurious fixed points).
+There are no fixed points $(U,D)$ of the algorithm in (3) such that $\|\bar{U}\|_{F}>\alpha$.
+As a consequence, all valid fixed points satisfy $\|\bar{U}\|_{F}\leq\alpha$, and the fixed point structure coincides with that of BM factorization (after the change of variables $U\mapsto UD^{1/2}$).
+Thus, adding the diagonal factor $D$ and bounding $U$ does not create new fixed points for $X$.
+Proof.
+Suppose $\|\bar{U}\|_{F}>\alpha$ and $\lambda_{j}>0$. Then, by combining Lemma 1(b)\,(b) and $(d)$, we get
+$0\overset{(d)}{=}u_{j}^{\top}\nabla f(X)u_{j}\overset{(b)}{=}-\tfrac{\beta}{\lambda_{j}}\|u_{j}\|^{2}.$
+which forces $u_{j}=0$.
+If instead $\lambda_{j}=0$, then $(b)$ again implies $u_{j}=0$.
+Thus in both cases the columns of $U$ must vanish, leading to $U=0$, and hence $\bar{U}=0$, which contradicts the assumption $\|\bar{U}\|_{F}>\alpha$.
+Hence, for any fixed point of the algorithm, the constraint is inactive and $\|\bar{U}\|_{F}\leq\alpha$.
+It remains to characterize these fixed points.
+From Lemma 1(a)(a), any fixed point satisfies $\nabla f(X)u_{j}\lambda_{j}=0$ for all $j$.
+Define $\hat{u}_{j}:=u_{j}\sqrt{\lambda_{j}}$ and let $\hat{U}=[\hat{u}_{1},\ldots,\hat{u}_{r}]$.
+Then $\hat{U}\hat{U}^{\top}=UDU^{\top}=X$ and $\nabla f(X)\hat{u}_{j}=0$ for all $j$, which coincides with the fixed-point condition of the BM factorization.
+Hence, the fixed points of our method correspond exactly to those of BM.
+Conversely, let $\hat{U}$ be a stationary point of the BM factorization, so $X=\hat{U}\hat{U}^{\top}$ and $\nabla f(X)\hat{U}=0$.
+For any $\alpha>0$, set $U=(\alpha/\|\hat{U}\|_{F})\hat{U}$ and $D=(\|\hat{U}\|_{F}^{2}/\alpha^{2})I$.
+Then $\|U\|_{F}=\alpha$, $UDU^{\top}=X$, and $\nabla f(X)UD=0$.
+Thus, every stationary point of BM corresponds to a stationary point of UDU, and vice versa, at the level of $X=UDU^{\top}=\hat{U}\hat{U}^{\top}$.
+∎
+Although the fixed points of UDU coincide with those of BM, the above analysis still provides insight into the low-rank bias of the algorithm.
+In particular, when $U$ grows such that $\lVert\bar{U}\rVert\geq\alpha$, the updates tend to amplify directions aligned with the negative eigenvectors of $\nabla f(X)$.
+Concretely, we can express the update rules in terms of the columns of $u_{j}$ and $\lambda_{j}$ as $\bar{u}_{j}=u_{j}-2\eta\nabla f(X)u_{j}\lambda_{i}$ and $\smash{\bar{\lambda}_{j}=\lambda_{j}-\eta u_{j}^{\top}\nabla f(X)u_{j}}$.
+These expressions show that both $\bar{u}_{j}$ and $\bar{\lambda}_{j}$ tend to grow in the directions aligned with the negative eigenvectors of $\nabla f(X)$, even though, by Lemma 1(d)(d), no fixed point with $\lambda_{j}>0$ is possible unless $u_{j}^{\top}\nabla f(X)u_{j}=0$.
+This tension suggests that the dynamics either suppress some columns of $U$ (via projection) or drive $X$ toward configurations where $\nabla f(X)u_{j}=0$, thereby reinforcing low-rank structure along the path to convergence.
+It is worth noting, that this discussion is a simplified perspective aimed at gaining intuition.
+In reality, the alignment or shrinking of the columns of $U$ and the minimization of $f(X)$ along these directions occur simultaneously and interact in a complex manner.
+Nevertheless, we can clearly observe these effects in our numerical experiments.
+In Section A.3, we present the evolution of the column norms of $U$ and the diagonal entries of $D$ over the iterations in our matrix completion experiment.
+Our results show that initially, a few specific columns of $U$ grow, pushing the others numerically to zero.
+Once $f(X)$ is effectively minimized with respect to these initial columns, additional columns are identified and begin to grow.
+Eventually, the algorithm converges to a low-rank solution, with $UDU^{\top}$ becoming rank-revealing since only a few columns of $U$ remain nonzero.
+We further observe that these nonzero columns are orthogonal, highlighting how the algorithm’s tendency to align $u_{j}$ with negative eigenvectors of $\nabla f(X)$ implicitly induces structured low-rank solutions along the path.
+3 Neural Networks with Diagonal Hidden Layers
+Figure 3: UDV structure. The weights in diagonal layer $D$ are denoted as $w_{j}$.
 This section extends our approach to neural networks.
-Consider a dataset comprising $n$ data points $(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathbb{R}^{d}\times\mathbb{R}^{c}$.
-We first define a three-layer neural network defined as
-$\phi(\mathbf{x}):=\sum_{j=1}^{m}\mathbf{v}_{j}w_{j}\mathbf{u}_{j}^{\top}\mathbf{x}~{}~{}\approx~{}~{}\mathbf{y}.$ (6)
-The first and third layers are fully connected, and the middle is a diagonal layer, as illustrated in Figure 2.
+We introduce a three-layer neural network, with the first and third layers fully connected and the middle layer diagonal, as illustrated in Figure 3.
+Let $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}\subset\mathbb{R}^{d}\times\mathbb{R}^{c}$ denote a dataset of size $n$. The resulting network is given by
+$\phi(\mathbf{x}):=\sum\nolimits_{j=1}^{m}\mathbf{v}_{j}w_{j}\mathbf{u}_{j}^{\top}\mathbf{x}$ (4)
+with the goal of approximating $\phi(\mathbf{x}_{i})\approx\mathbf{y}_{i}$.
 Drawing parallels between our matrix factorization model in (2) and neural network training, we impose Euclidean norm constraints on the weights of the fully connected layers.
 Under these conditions, the training problem can be formulated as follows:
-$\min_{\begin{subarray}{c}\mathbf{u}_{j}\in\mathbb{R}^{d}\\ \mathbf{v}_{j}\in\mathbb{R}^{c}\\ w_{j}\in\mathbb{R}\end{subarray}}$ $\frac{1}{2n}\sum_{i=1}^{n}\|\sum_{j=1}^{m}\mathbf{v}_{j}\,w_{j}\,\mathbf{u}_{j}^{\top}\mathbf{x}_{i}-\mathbf{y}_{i}\|_{2}^{2}$ (7)
-$\mathrm{subj.to}$ $\sum_{j=1}^{m}\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~{}~{}\sum_{j=1}^{m}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1,~{}~{}\text{and }~{}~{}w_{j}\geq 0;\quad\text{for all}~{}~{}j=1,\ldots,m.$
-The norm constraints in our training problem can be interpreted as a stronger form of weight decay, one of the most commonly used regularization techniques in neural networks, which lends further justification to our formulation.
+$\min_{\mathbf{u}_{j},w_{j},\mathbf{v}_{j}}$ $\frac{1}{2n}\sum\nolimits_{i=1}^{n}\|\sum\nolimits_{j=1}^{m}\mathbf{v}_{j}\,w_{j}\,\mathbf{u}_{j}^{\top}\mathbf{x}_{i}-\mathbf{y}_{i}\|_{2}^{2}$ (5)
+$~\mathrm{s.t.}$ $\sum\nolimits_{j=1}^{m}\|\mathbf{u}_{j}\|_{2}^{2}\leq 1,~~\sum\nolimits_{j=1}^{m}\|\mathbf{v}_{j}\|_{2}^{2}\leq 1,$
+$\text{and }~~w_{j}\geq 0;\quad\text{for all}~~j=1,\ldots,m.$
+The norm constraints can be interpreted as a stronger form of weight decay, a standard regularization in neural networks, which lends further justification to our formulation.
 We refer to this neural network structure as UDV.
-Figure 2: UDV structure. The weights in diagonal layer $D$ are denoted as $w_{j}$.
-4.1 Numerical Experiments on Neural Networks
-In this section, we test the proposed UDV framework on regression and classification tasks, comparing it with fully connected two-layer neural networks (denoted as UV in the subsequent text) using both linear and ReLU activation functions.
-This comparison is fair in terms of computational cost, as the cost incurred by the diagonal layer —which can also be viewed as a parameterized linear activation function— is negligible.
-We demonstrate a strong empirical bias toward low-rank solutions in all our experiments.
+3.1 Experiments on Neural Networks
+In this section, we test the UDV structure on regression and classification tasks, comparing it with fully connected two-layer neural networks (denoted as UV in the subsequent text).
+This comparison is fair in terms of computational cost since the overhead introduced by the diagonal layer, which can also be regarded as a parameterized linear activation function, is negligible.
+To further ensure fairness in comparison, no additional fine-tuning was introduced for the proposed method.
+We observe a strong empirical bias toward low-rank solutions in all our experiments with the UDV structure.
 We also present a proof-of-concept use case of this strong bias, combined with an SVD-based pruning strategy, to produce compact networks.
 3.1.1 Implementation Details
-Computing environment.
-All classification tasks were conducted on an NVIDIA A100 GPU with four cores of the AMD Epyc 7742 processor, while regression tasks were conducted on a single core of an Intel Xeon Gold 6132 processor.
-We used Python 3.9.5 and PyTorch 2.0.1.
 Datasets.
 We used two datasets for the regression tasks: House Prices - Advanced Regression Techniques (HPART) (Anna Montoya, 2016) and New York City Taxi Trip Duration (NYCTTD) (Risdal, 2017).
-We allocated 80% of the data for the training and reserved the remaining 20% for validation.
+We allocated 80% of the data for the training and reserved the remaining 20% for test.
 Following (Huang, 2003), we set the number of hidden neurons in the diagonal layer $\smash{m=\texttt{round}\big(\sqrt{(c+2)d}+2\sqrt{d/(c+2)}\big)}$.
 This results in a network structure ($d$-$m$-$c$) of 79-26-1 for HPART, and 12-10-1 for NYCTTD.
-For classification tasks, we used the normalized MNIST dataset (LeCun et al., 2010).
-We applied transfer learning by replacing the classifier layers of three advanced neural networks –MaxViT-T (Tu et al., 2022), EfficientNet-B0 (Tan & Le, 2019), and RegNetX-32GF (Radosavovic et al., 2020)– with UDV, while using pre-trained weights from ImageNet-1K (Deng et al., 2009).
-Specifically, we retained all layers up to the first fully connected layer of the classifier and replaced the subsequent layers.
-The number of hidden neurons in the diagonal layer was set to as $m=\texttt{floor}(\tfrac{2}{3}d)$.
-This results in a UDV network structure ($d$-$m$-$c$) of 512-341-10 for MaxViT-T, 1280-853-10 for EfficientNet-B0, and 2520-1680-10 for RegNetX-32GF.
+For classification tasks, we used the CIFAR-10 and CIFAR-100 dataset (Krizhevsky, 2009).
+We applied transfer learning to three modern high-performing architectures: MaxViT-T (Tu et al., 2022), EfficientNet-B0 (Tan & Le, 2019), and RegNetX-32GF (Radosavovic et al., 2020).
+The original classification head (the final linear layer) of each model was replaced with two linear layers, forming the backbone classifier (UV).
+By inserting a diagonal layer between them, the classifier was transformed into the proposed UDV structure.
+The new linear layers were initialized with a truncated normal distribution, and the diagonal layer was initialized as an identity matrix.
+The vector dimension followed that of the preceding layer, resulting in a UDV classifier structure ($d$-$m$-$c$), where $c$ denotes the number of classes (10 for CIFAR-10 and 100 for CIFAR-100).
+Specifically, the structures are 512-512-$c$ for MaxViT-T, 1280-1280-$c$ for EfficientNet-B0, and 2520-2520-$c$ for RegNetX-32GF.
 Loss function. We used mean squared error for regression and cross-entropy loss for classification.
@@
 We tested the results using four different optimization algorithms for training: Adam (Kingma & Ba, 2014), Mini-Batch Gradient Descent (MBGD) (LeCun et al., 2002), NAdam (Dozat, 2016), and Mini-Batch Gradient Descent with Momentum (MBGDM) (Sutskever et al., 2013).
-For classification, we used different batch sizes for different models: 128 for MaxViT-T and RegNetX-32GF, and 384 for EfficientNet-B0.
-We tuned the step size for all models and optimization algorithms: For regression tasks, we tested step sizes $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$, $2$, $3$.
-Larger step sizes ($1$, $2$, $3$) were often excluded for the UV model due to divergence.
-In classification experiments, we tested learning rates $10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$ with Adam and NAdam; and we tested $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$, $2$, $3$, $5$ with MBGD and MBGDM.
-Training Procedure.
-UV and UDV models were initialized identically, with the diagonal elements of $D$ initialized using Kaiming Uniform Initialization, consistent with the default initialization for fully connected layers in PyTorch.
-The models were trained for 200 epochs on HPART, 50 epochs on NYCTTD, and 70 epochs on MNIST; and results were averaged over 1000 random seeds for HPART, 100 for NYCTTD, and 1 for MNIST to ensure robustness.
-The validation loss, used as a generalization metric in regression, was averaged over the final 20 epochs for the HPART dataset and the final 5 epochs for the NYCTTD dataset.
-Similarly, validation accuracy for classification tasks was averaged over the last 5 epochs to ensure stability in the reported values.
-Repository. The code is available at https://github.com/Titanium-H/UDV
-4.1.2 Low-rank Bias in Neural Network Training
-Table 1 presents the validation loss (for regression) or validation accuracy (for classification) of the UDV model compared to the classical UV model with linear and ReLU activation functions.
+We tuned the learning rates (LRs, the term learning rates is used instead of step sizes for clarity in the machine learning context) for all models and optimization algorithms.
+In addition, a cosine annealing learning rate scheduler was applied to stabilize training and improve convergence.
+Training Procedure and Metrics.
+The models were trained for 200 epochs on HPART, 50 epochs on NYCTTD, and 100 epochs on CIFAR-10 and CIFAR-100.
+Test loss for regression was averaged over the last 20 and 5 epochs, respectively, whereas test accuracy for classification was averaged over the final 3 epochs.
+3.1.2 Low-rank Bias in Neural Networks
+Table 1 presents the test loss (for regression) or test accuracy (for classification) of the UDV model compared to the classical UV model.
 For each configuration (dataset and model architecture), the results are obtained by selecting the best algorithm and learning rate pair.
-Moreover, Figure 3 illustrates the singular value spectrum of the solutions corresponding to each entry in these tables.
-We focus on the singular values from the $U$ and $UD$ layers, as they generate the primary data representation, while omitting the $V$ layer, which serves as the feature selection layer and is a tall matrix by definition, given that $c\ll m$ in most cases.
-Collectively, these results show that the UDV framework achieves competitive prediction accuracy while exhibiting a strong implicit bias toward low-rank solutions, as indicated by the faster decay in the singular value spectrum.
-Table 1: Model performance using different models. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
-Tasks Regression (Test Loss) Classification (Test Accuracy)
-Dataset HPART NYCTTD MNIST
-UDV
-$1.304\times 10^{-3}$
-Adam: $10^{-3}$
-$5.248\times 10^{-6}$
-NAdam: $10^{-4}$
-M: 99.67%
-MBGDM: $10^{-2}$
-E: 99.63%
-MBGDM: $10^{-1}$
-R: 99.74%
-MBGDM: $10^{-2}$
-UV
-$1.333\times 10^{-3}$
-Adam: $10^{-3}$
-$5.251\times 10^{-6}$
-Adam: $10^{-3}$
-M: 99.69%
-MBGDM: $10^{-2}$
-E: 99.60%
-Adam: $10^{-3}$
-R: 99.66%
-MBGD: $10^{0}$
-UV-ReLU
-$1.167\times 10^{-3}$
-Adam: $10^{-3}$
-$5.323\times 10^{-6}$
-NAdam: $10^{-3}$
-M: 99.68%
-NAdam: $10^{-4}$
-E: 99.68%
-MBGDM: $10^{-1}$
-R: 99.73%
-MBGD: $10^{0}$
-Figure 3: Singular value spectrum corresponding to the solutions reported in Table 1.
-4.1.3 Reducing Network Size with SVD-based Pruning
+On CIFAR-100, the UDV model underperforms UV by about 1% in test accuracy.
+We note, however, that UDV achieves its best performance earlier in training (around 90 epochs), suggesting that early stopping may close this gap.
+Table 1: Comparison of models with their best performance. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
+Tasks Regression (Test Loss) Classification (Test Accuracy %)
+Dataset HPART NYCTTD CIFAR-10 CIFAR-100
+UDV $1.304\times 10^{-3}$ $5.248\times 10^{-6}$
+M: 99.17
+E: 97.87
+R: 98.67
+M: 86.18
+E: 91.63
+R: 89.79
+UV $1.333\times 10^{-3}$ $5.251\times 10^{-6}$
+M: 99.11
+E: 97.99
+R: 98.65
+M: 87.75
+E: 92.03
+R: 90.33
+Figure 5 (top row) shows the singular value spectra for RegNetX-32GF on CIFAR-100 under different optimization algorithms.
+We report the spectra for the $U$ (for the classical model) and $UD$ (for the proposed model) layers, which generate the main data representation, and omit the V layers since it primarily acts as a feature selector and is typically tall ($c\ll m$).
+Collectively, the results show that the UDV framework matches the predictive accuracy of UV while exhibiting a clear implicit bias toward low-rank solutions, as reflected in the faster spectral decay.
+Figure 4 presents similar results for the HPART regression task. The findings are consistent, as UDV induces a strong low-rank bias while maintaining predictive performance comparable to UV.
+Figure 4: Singular value spectra of the solutions obtained with UV and UDV formulations on the HPART regression task under different optimization algorithms.
+3.1.3 Reducing Network Size via SVD-based Pruning
 Efficient and lightweight feed-forward layers are crucial for real-world applications.
 For instance, the Apple Intelligence Foundation Models (Gunter et al., 2024) recently reported that pruning hidden dimensions in feed-forward layers yields the most significant gains in their foundation models.
-Building on this insight, we leverage the inherent low-rank bias of the UDV architecture through an SVD-based pruning strategy to produce compact networks without sacrificing performance.
+Building on this insight, we leverage the low-rank bias of the UDV architecture through an SVD-based pruning strategy to produce compact networks without sacrificing performance.
 A low-rank solution was observed when applying SVD to $UD$ layers:
-$UD=\texttt{USV}^{\top},\quad\texttt{U}\in\mathbb{R}^{d\times m},\quad\texttt{S}\in\mathbb{R}^{m\times m},\quad\texttt{V}^{\top}\in\mathbb{R}^{m\times m}.$ (8)
+$UD=\texttt{USV}^{\top},\quad\texttt{U}\in\mathbb{R}^{d\times m},\quad\texttt{S}\in\mathbb{R}^{m\times m},\quad\texttt{V}^{\top}\in\mathbb{R}^{m\times m}.$
 By dropping small singular values in S, these matrices can be truncated to $\smash{\bar{\texttt{U}}\in\mathbb{R}^{d\times r}}$, $\smash{\bar{\texttt{S}}\in\mathbb{R}^{r\times r}}$ and $\bar{\texttt{V}}^{\top}\in\mathbb{R}^{r\times m}$, where $0<r<m$.
-Consequently, ($m-r$) neurons can be pruned, and new weight matrices are assigned:
-$\bar{U}=\bar{\texttt{U}}\in\mathbb{R}^{d\times r},\quad\bar{D}=\bar{\texttt{S}}\in\mathbb{R}^{r\times r},\quad\bar{V}=\bar{\texttt{V}}^{T}V\in\mathbb{R}^{r\times c}.$ (9)
-We applied this pruning strategy on the models from Table 1.
-The left part of Figure 4 presents an example comparing the generalization capability of pruned models.
-For comparison, we created compact models by training from scratch with a reduced number of neurons $m$ in the hidden layer.
-The performance change for these models is shown in the right panel of Figure 4.
-Although our pruned networks derived from UDV demonstrate that models with significantly fewer parameters can still achieve strong generalization, these compressed architectures are often more challenging to optimize directly within the reduced space, consistent with prior findings in the literature (Arora et al., 2018; Chang et al., 2021).
-We omit the results for retraining with the UV model, as they show similar trends to UDV in this context, though UDV generally exhibits superior generalization.
-Figure 4: Comparison of SVD-based pruning and re-training compact networks on the HPART dataset using the NAdam algorithm with a learning rate of $10^{-3}$.
-Negative percentages indicate improvements over the baseline.
-SVD-based pruning demonstrates that the UDV leads to a compact model without performance degradation, while retraining shows that the UDV achieves better generalization in a compact model.
+Consequently, (mrm-r) neurons can be pruned, and new weight matrices are assigned:
+$\bar{U}=\bar{\texttt{U}}\in\mathbb{R}^{d\times r},\quad\bar{D}=\bar{\texttt{S}}\in\mathbb{R}^{r\times r},\quad\bar{V}=\bar{\texttt{V}}^{\top}V\in\mathbb{R}^{r\times c}.$
+The bottom row of Figure 5 shows test accuracy of the post-training pruned models as a function of the number of remaining neurons.
+The x-axis is cropped for clarity (the full model has 2520 neurons before pruning).
+The UDV-based models maintain high accuracy even under aggressive pruning and consistently outperform the UV baseline at comparable compression levels.
+This demonstrates that UDV induces low-rank representations that are more amenable to compression without compromising performance.
+Although the pruned networks show that models with significantly fewer parameters can still generalize well, training such small models from scratch typically results in lower test accuracy, consistent with prior findings in the literature (Arora et al., 2018; Chang et al., 2021).
+Figure 5: Singular value spectra and post-training SVD-based pruning results for RegNetX-32GF on CIFAR-100 under different optimization algorithms.
+[Top] UDV induces faster spectral decay than the baseline UV model.
+[Bottom] Test accuracy as a function of remaining neurons after pruning (full model has 2520 neurons, x-axis cropped for clarity).
+UDV-based networks retain high accuracy under aggressive pruning
 3.1.4 Further Details and Discussions
@@
 Despite these differences, the UDV architecture consistently demonstrated a strong bias toward low-rank solutions.
-Additional experiments and details can be found in the Appendices, with key results summarized below:
-  • In the early stages of this research, we explored four variants of UDV, each differing slightly in their constraints.
-We selected the version presented in (7), as it generally exhibits the most pronounced decay in the singular value spectrum.
-For completeness, details of the other three variants are provided in Section B.1.
-  • Section B.2 provides additional results for the experiment described in Section 4.1.2.
-Specifically, we present results analogous to those in Table 1 and Figure 3, but focusing exclusively on the MBGDM algorithm.
-These results exhibit similar trends, reinforcing consistency across different methods.
-Additionally, comprehensive performance comparisons across all algorithms and models are provided in Tables SM2, SM3, SM4 and SM5 in the Appendices.
-  • Additional results on SVD-based pruning are provided in Section B.3.
-We demonstrate that the UDV framework consistently achieves low-rank solutions across various problem configurations.
-Furthermore, we analyze the effect of learning rate on the singular value spectrum, similar to the analysis in Figure 1, but applied to neural network experiments.
+Additional experiments can be found in the Appendices, with key results summarized below.
+Section B.1 provides comprehensive results complementing those in Sections 3.1.2 and 3.1.3, covering other configurations (MaxViT-T and EfficientNet-B0) and datasets (CIFAR-10/100).
+The low-rank patterns persist across settings, and the performance gap in pruning between UDV and UV becomes more pronounced as task complexity increases, highlighting the growing advantage of low-rank structure.
+Early in this work we explored four UDV variants with different constraints.
+All produced low-rank solutions, but the formulation in (5) showed the strongest spectral decay, and we adopt it as the main model.
+The remaining variants are described in Section B.2.
+Additional experiments on SVD-based pruning with these UDV variants are provided in Section B.3.
+Furthermore, we analyzed the effect of learning rate on the singular value spectrum, similar to the analysis in Figure 1, but applied to neural network experiments.
 This analysis confirms that the UDV framework produces low-rank solutions across a broad range of learning rates.
-  • Section B.4 extends the UDV framework by incorporating ReLU activation.
-Preliminary experiments with the UDV-ReLU model indicate that it also tends to yield low-rank solutions, similar to those observed in the original UDV framework.
-  • Prior work on implicit bias in neural networks suggests that increasing depth enhances the tendency toward low-rank solutions (Arora et al., 2019; Feng et al., 2022), raising the question of whether the pronounced bias in the UDV framework is just a consequence of adding a diagonal layer.
-To investigate this, we conducted experiments comparing the UDV model to fully connected three-layer networks, as detailed in Section B.5.
-Additionally, we included a UDV model without constraints in these comparisons.
-The results indicate that this bias cannot be attributed solely to depth, highlighting the critical role of explicit constraints.
+Section B.4 extends the UDV framework by incorporating ReLU activation. Preliminary experiments suggest that UDV-ReLU also promotes low-rank solutions, consistent with the original UDV model.
+Prior work on implicit bias in neural networks suggests that network depth can enhances the tendency toward low-rank solutions (Arora et al., 2019; Feng et al., 2022), raising the question of whether the pronounced bias in UDV is just a consequence of the extra diagonal layer.
+To test this, we compared UDV with fully connected three-layer networks in Section B.5.
+We also examined an unconstrained UDV variant.
+The results show that the bias cannot be explained by depth alone, highlighting the importance of explicit constraints.
 Section B.6 compares the spectral decay of the UDV network to that induced by classical weight decay regularization in two- and three-layer networks.
-The results show that while weight decay promotes singular value decay, it can not reproduce the strong decay observed in the UDV model.
+While weight decay promotes singular value decay, it can not reproduce the strong decay observed in the UDV model.
+Finally, Section B.7 presents a toy example applying the UDV structure within the LoRA framework to fine-tune a pre-trained LLaMA-2 model on a causal language modeling task.
+As a proof of concept, the pruning results suggest that UDV blocks could serve as drop-in replacements for linear layers, showing promising performance under compression.
+Most importantly, our comprehensive set of experiments shows that the UDV framework consistently induces a strong implicit bias toward low-rank solutions across diverse architectures, datasets, training setups and algorithms.
 4 Conclusions
-We proposed a new matrix factorization framework, inspired by the observation that implicit bias is driven by dynamics—potentially divergent—that are distinct from those leading to convergence of the objective function.
-This framework constrains the factors within Euclidean norm balls and introduces a middle diagonal factor to ensure the search space is not restricted.
-Numerical experiments demonstrate that this approach significantly strengthens the low-rank bias in the solution.
+We introduced a new matrix factorization framework that constrains the factors within Euclidean norm balls and incorporates a diagonal middle factor to avoid restricting the search space.
+Numerical experiments show that this formulation enhances the implicit bias toward low-rank solutions.
 To explore the broader applicability of our findings, we designed an analogous neural network architecture with three layers, constraining the fully connected layers and adding a diagonal hidden layer, referred to as UDV.
-Extensive experiments show that the proposed UDV architecture achieves competitive performance compared to standard fully connected networks, while inducing a structured solution with a strong bias toward low-rank representations.
-Additionally, we explored the utility of this low-rank structure by applying an SVD-based pruning strategy, illustrating how it can be leveraged to construct compact networks that are more efficient for downstream tasks.
+Extensive experiments (including those in Appendices) show that the UDV architecture induces a strong bias toward low-rank representations.
+Additionally, we demonstrated the utility of this low-rank structure by applying an SVD-based pruning strategy, as a proof of concept, illustrating how it can be leveraged to construct compact networks that are more efficient for downstream tasks.
+The proposed model exhibits reduced rank regression behavior, where the training process gradually increases the model rank.
+We believe these our should be of broader interest to research on implicit bias.
+Although we provide some theoretical insights, developing a more complete theory and designing algorithms that fully exploit the model’s regularization capabilities remain important directions for future work.
 Acknowledgments
@@
 [Right] Singular value spectrum after $10^{6}$ iterations.
-A.2 UDU Factorization Produces Rank-Revealing Solutions
+A.2 Phase Retrieval
+A.2.1 Phase Retrieval on Synthetic Data
+The experiment with real data in Section 2.2.2 is large-scale and thus impractical for testing full-dimensional factorizations with $r=d$.
+To complement it, we perform a small-scale synthetic test bed for phase retrieval.
+We consider the matrix sensing formulation of phase retrieval (Candes et al., 2013), where the goal is to recover a signal from quadratic measurements of the form $y_{i}=|\langle a_{i},x\rangle|^{2}$.
+Although the standard maximum likelihood estimators lead to a non-convex optimization problem due to the quadratic terms, the problem can be reformulated as minimizing a convex function under a rank constraint in a lifted space.
+By denoting the lifted variable as $X=xx^{\top}$, the measurements can be expressed as
+$\smash{y_{i}=\langle a_{i}^{\top}x,a_{i}^{\top}x\rangle=\langle a_{i}a_{i}^{\top},xx^{\top}\rangle:=\langle A_{i},X\rangle,}$
+which allows us to reformulate the problem as
+$\smash{\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{s.t.}\quad\mathrm{rank}(X)\leq 1.}$
+In this experiment, we use a gray-scale image of size $16\times 16$ pixels, selected from the Pixel Art dataset (Elgazar, 2024), as the signal $x\in\mathbb{R}^{n}$ to recover, where $n=256$.
+We generate a synthetic measurement system by sampling $a_{1},\ldots,a_{m}$ from the standard Gaussian distribution, with $m=2n$.
+We then solve problems (1) and (2).
+We initialized $U_{0}\in\mathbb{R}^{n\times n}$ with entries drawn iid from the standard Gaussian distribution, then rescaled to have a unit Frobenius norm.
+For $D_{0}\in\mathbb{R}^{n\times n}$, we used the identity matrix.
+We used step-size $\eta=\frac{1}{L}$ where $L$ denotes the smoothness constant.
+After solving the problem, we recover $x\in\mathbb{R}^{n}$ from the lifted variable $X\in\mathbb{R}^{n\times n}$ by selecting its dominant eigenvector.
+Figure SM2 demonstrates that the proposed method consistently identifies low-rank solutions, in line with the results of our other experiments.
+Figure SM3 displays the recovered image, demonstrating that the proposed method achieves higher quality than BM factorization within the same number of iterations, which can be attributed to the low-rank structure of the solution.
+Figure SM2: Comparison of singular value spectrum in recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
+Figure SM3: Comparison of recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
+A.2.2 Intermediate Reconstructions in Fourier Ptychography Imaging
+We provide additional details for the Fourier ptychography imaging experiment in Section 2.2.2, showing the reconstructions at intermediate iterations (Figure SM4).
+$k=0$
+$k=10$
+$k=100$
+$k=1000$
+$k=10000$
+Figure SM4: Reconstructed images obtained from the proposed UDU factorization (top row) and the standard BM factorization (bottom row) at different iterations, starting from the same random initialization.
+While the UDU method converges to a clean and accurate recovery, the BM reconstructions display severe artifacts.
+A.3 UDU Factorization Produces Rank-Revealing Solutions
 The proposed factorization method naturally produces rank-revealing solutions.
@@
 This behavior resembles the mechanism of the power method and provides insights into its connection with divergent forces.
-In Figure SM2, we illustrate the evolution of the column norms of $U$ from the matrix completion experiment described in Section 3.2.
-Figure SM3 displays the corresponding entries in the diagonal factor $D$.
-For comparison, Figure SM4 shows the column norms of $U$ obtained from the same experiment using the standard BM factorization.
+In Figure SM5, we illustrate the evolution of the column norms of $U$ from the matrix completion experiment described in Section 2.2.
+Figure SM6 displays the corresponding entries in the diagonal factor $D$.
+For comparison, Figure SM7 shows the column norms of $U$ obtained from the same experiment using the standard BM factorization.
 The results are shown for $\xi=10^{-1}$ and $\eta=10^{-1}$ over $10^{5}$ iterations, but the findings are similar across other parameter settings.
 We also repeat the experiment with the noisy measurements described in Section A.1.
-The results are shown in Figures SM5, SM6 and SM7.
-Figure SM2: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization.
+The results are shown in Figures SM8, SM9 and SM10.
+Figure SM5: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization.
 The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
-Figure SM3: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization. The x-axis represents indices.
-Figure SM4: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization, shown for comparison.
+Figure SM6: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization. The x-axis represents indices.
+Figure SM7: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization, shown for comparison.
 The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
-Figure SM5: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization with noisy measurements.
+Figure SM8: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization with noisy measurements.
 The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
-Figure SM6: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents indices.
-Figure SM7: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization with noisy measurements, shown for comparison.
+Figure SM9: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents indices.
+Figure SM10: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization with noisy measurements, shown for comparison.
 The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
-A.3 Phase Retrieval Image Recovery
-We conduct a numerical experiment on the matrix sensing problem arising in phase retrieval image recovery (Candes et al., 2013).
-The objective is to recover a signal from its quadratic measurements of the form $y_{i}=|\langle a_{i},x\rangle|^{2}$.
-Although the standard maximum likelihood estimators lead to a non-convex optimization problem due to the quadratic terms, the problem can be reformulated as minimizing a convex function under a rank constraint in a lifted space.
-By denoting the lifted variable as $X=xx^{\top}$, the measurements can be expressed as
-$\smash{y_{i}=\langle a_{i}^{\top}x,a_{i}^{\top}x\rangle=\langle a_{i}a_{i}^{\top},xx^{\top}\rangle:=\langle A_{i},X\rangle,}$
-which allows us to reformulate the problem as
-$\smash{\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{subj.\ to}\quad\mathrm{rank}(X)\leq 1,}$
-representing a special case of problem (1).
-In this experiment, we use a gray-scale image of size $16\times 16$ pixels, selected from the Pixel Art dataset (Elgazar, 2024), as the signal $x\in\mathbb{R}^{n}$ to recover, where $n=256$.
-We generate a synthetic measurement system by sampling $a_{1},\ldots,a_{m}$ from the standard Gaussian distribution, with $m=2n$.
-We then solve problems (2) and (4).
-We initialized $U_{0}\in R^{n\times n}$ with entries drawn iid from the standard Gaussian distribution, then rescaled to have a unit Frobenius norm.
-For $D_{0}\in R^{n\times n}$, we used the identity matrix.
-We used step-size $\eta=\frac{1}{L}$ where $L$ denotes the smoothness constant.
-After solving the problem, we recover $x\in\mathbb{R}^{n}$ from the lifted variable $X\in\mathbb{R}^{n\times n}$ by selecting its dominant eigenvector.
-Figure SM8 demonstrates that the proposed method consistently identifies low-rank solutions, in line with the results of our other experiments.
-Figure SM9 displays the recovered image, demonstrating that the proposed method achieves higher quality than BM factorization within the same number of iterations, which can be attributed to the low-rank structure of the solution.
-Figure SM8: Comparison of singular value spectrum in recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
-Figure SM9: Comparison of recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
+A.4 Additional Details on the Fixed Point Analysis
+In Section 2.3, we provided insights into the inner workings of the UDU framework on matrix factorization problems based on a fixed-point analysis. Recall the algorithm steps:
+$U_{k+1}$ $=\Pi_{U}(\bar{U}_{k+1})$ $\text{where}\quad\bar{U}_{k+1}=U_{k}-2\eta\nabla f(U_{k}D_{k}U_{k}^{\top})U_{k}D_{k}$
+$D_{k+1}$ $=\Pi_{D}(\bar{D}_{k+1})$ $\text{where}\quad\bar{D}_{k+1}=D_{k}-\eta U_{k}^{\top}\nabla f(U_{k}D_{k}U_{k}^{\top})U_{k}$
+Let $(U,D)$ denote a fixed point of this algorithm and let $X=UDU^{\top}$. Suppose $\lVert\bar{U}\rVert\leq\alpha$. In this case, the projection step $\Pi_{U}$ acts as an identity:
+$U=\bar{U}=U-2\eta\nabla f(X)UD\quad\implies\quad\nabla f(X)UD=0$ (6)
+Otherwise, if we assume $\lVert\bar{U}\rVert\geq\alpha$, then the projection $\Pi_{U}$ has a normalization effect:
+$U=\frac{\alpha}{\lVert\bar{U}\rVert}\bar{U}=\frac{\alpha}{\lVert\bar{U}\rVert}\big(U-2\eta\nabla f(X)UD\big)$
+which we can reorganize as
+$\nabla f(X)UD=-\beta U\quad\text{where}\quad\beta:=\frac{\lVert\bar{U}\rVert-\alpha}{2\alpha\eta}.$ (7)
+Denoting the columns of $U$ by $u_{j}$ and the diagonal entries of $D$ by $\lambda_{j}$, we express the conditions (6) and (7) in terms of the individual components as presented in conditions $(a)$ and $(b)$ in Section 2.3.
+Next, we investigate the $D$ component.
+Note that the projection $\Pi_{D}$ operates in an entry-wise separable manner; it maps all non-diagonal entries to 0, and takes the positive part for the diagonal entries.
+It maps all off-diagonal entries to zero and applies the positive-part operator to the diagonal entries.
+Hence, it suffices to examine only the diagonal elements.
+Suppose $\bar{\lambda}_{j}\leq 0$. Then $\lambda_{j}=\max(\bar{\lambda}_{j},0)=0$, and
+$\bar{\lambda}_{j}=\lambda_{j}-\eta\,u_{j}^{\top}\nabla f(X)u_{j}=-\eta\,u_{j}^{\top}\nabla f(X)u_{j}\quad\implies\quad u_{j}^{\top}\nabla f(X)u_{j}\geq 0.$ (8)
+Otherwise, if $\bar{\lambda}_{j}>0$, then $\lambda_{j}=\max(\bar{\lambda}_{j},0)=\bar{\lambda}_{j}$, hence
+$\lambda_{j}=\bar{\lambda}_{j}=\lambda_{j}-\eta\,u_{j}^{\top}\nabla f(X)u_{j}\quad\implies\quad u_{j}^{\top}\nabla f(X)u_{j}=0.$ (9)
+The conditions in (8) and (9) correspond to cases $(c)$ and $(d)$ in Section 2.3.
 Appendix B Additional Details on Neural Network Experiments
-B.1 Design Variants of the UDV Architecture
+Computing environment.
+All classification tasks were conducted on an NVIDIA A100 GPU with four cores of the AMD Epyc 7742 processor, while regression tasks were conducted on a single core of an Intel Xeon Gold 6132 processor.
+We used Python 3.12.3 and PyTorch 2.5.1.
+Robustness Evaluation. All reported results in the tables were averaged across 1000 seeds on HPART, 100 on NYCTTD, and 10 on CIFAR-10, CIFAR-100, and MNIST.
+B.1 Full Results of Neural Networks Reported in the Main Text
+Here we provide all results related to the image classification experiments discussed in Section 3.
+Table SM1 and Table SM2 present the complete results corresponding to Table 1 in the main text, while Figures SM11, SM12, SM13, SM14, SM15 and SM16 show all singular value spectrum and pruning results.
+The regression experiments were part of the earlier stage of this study, and their details are reported in Section B.2.
+Meanwhile, the early experiments reduced the need for extensive hyperparameter selection in the current study.
+Therefore, for Adam and NAdam we used learning rates of $10^{-3}$ and $10^{-4}$, whereas for MBGD and MBGDM we used $10^{-1}$ and $10^{-2}$.
+Table SM1: Comparison of model performance on CIFAR-10. Note “–” denotes results that are not applicable.
+Model MaxVit-T [M] EfficientNet-B0 [E] RegNetX-32GF [R]
+Optimizer Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM
+LR: $10^{-4}$
+UDV:99.14
+UV:99.11
+UDV:99.17
+UV:99.11
+UDV:97.87
+UV:97.98
+UDV:97.86
+UV:97.99
+UDV:98.67
+UV:98.64
+UDV:98.55
+UV:98.56
+LR: $10^{-3}$
+UDV:98.61
+UV:98.76
+UDV:98.72
+UV:98.69
+UDV:97.61
+UV:97.77
+UDV:97.68
+UV:97.69
+UDV:97.84
+UV:97.60
+UDV:97.78
+UV:97.68
+LR: $10^{-2}$ - -
+UDV:98.04
+UV:98.25
+UDV:98.93
+UV:98.94
+UDV:95.98
+UV:96.65
+UDV:97.79
+UV:97.90
+UDV:98.00
+UV:98.18
+UDV:98.61
+UV:98.65
+LR: $10^{-1}$ - -
+UDV:98.93
+UV:99.01
+UDV:99.05
+UV:99.03
+UDV:97.86
+UV:97.82
+UDV:97.87
+UV:97.94
+UDV:98.67
+UV:98.64
+UDV:98.59
+UV:98.63
+Table SM2: Comparison of model performance on CIFAR-100. Note “–” denotes results that are not applicable and $\times$” denotes results that did not converge.
+Model MaxVit-T [M] EfficientNet-B0 [E] RegNetX-32GF [R]
+Optimizer Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM
+LR: $10^{-4}$
+UDV:84.69
+UV:87.75
+UDV:84.69
+UV:87.72
+UDV:91.58
+UV:91.90
+UDV:91.63
+UV:92.03
+UDV:89.17
+UV:89.22
+UDV:89.31
+UV:89.40
+LR: $10^{-3}$
+UDV:84.70
+UV:86.36
+UDV:84.76
+UV:86.09
+UDV:89.05
+UV:90.55
+UDV:89.49
+UV:90.60
+UDV:85.41
+UV:85.62
+UDV:85.73
+UV:85.44
+LR: $10^{-2}$ - -
+UDV: $\times$
+UV:84.57
+UDV:85.23
+UV:87.73
+UDV: $\times$
+UV:88.23
+UDV:89.20
+UV:91.55
+UDV: $\times$
+UV:88.82
+UDV:89.75
+UV:90.33
+LR: $10^{-1}$ - -
+UDV:85.45
+UV:87.74
+UDV:86.18
+UV:86.67
+UDV:89.88
+UV:91.56
+UDV:91.35
+UV:91.70
+UDV:89.79
+UV:90.30
+UDV:88.53
+UV:88.69
+Figure SM11: Singular value spectrum and pruning performance of MaxVit-T on CIFAR-10.
+Figure SM12: Singular value spectrum and pruning performance of EfficientNet-B0 on CIFAR-10.
+Figure SM13: Singular value spectrum and pruning performance of RegNetX-32GF on CIFAR-10.
+Figure SM14: Singular value spectrum and pruning performance of MaxVit-T on CIFAR-100.
+The learning rate of $10^{-2}$ for MBGD was unsuitable for our model, but no other hyperparameters were fine-tuned to ensure fairness.
+Figure SM15: Singular value spectrum and pruning performance of EfficientNet-B0 on CIFAR-100.
+The learning rate of $10^{-2}$ for MBGD was unsuitable for our model, but no other hyperparameters were fine-tuned to ensure fairness.
+Figure SM16: Singular value spectrum and pruning performance of RegNetX-32GF on CIFAR-100.
+The learning rate of $10^{-2}$ for MBGD was unsuitable for our model, but no other hyperparameters were fine-tuned to ensure fairness.
+B.2 Design Variants of the UDV Architecture
 When designing our network architecture, we considered four variants, including the one presented in Section 3. The three additional models were defined by the following sets of constraints:
@@
 Finally, UDV-v2 is identical to UDV-v1 but without the non-negativity constraints on the diagonal layer.
-We only report results from UDV, as defined in equation (7), in the main text.
-However, we provide results for the other variants in the Appendices for completeness and to illustrate the impact of different constraint settings on model performance.
-Figure SM11 and Figure SM12 show that UDV consistently finds a low-rank solution.
+Early image classification experiments were performed on the normalized MNIST dataset (LeCun et al., 2010), which allowed rapid iteration and efficient validation of the proposed method.
+The replacement of the classifier follows the procedure described in Section 3.1.
+The number of hidden neurons in the diagonal layer was defined as $m=\texttt{floor}(\tfrac{2}{3}d)$,and this results in a UDV network structure ($d$-$m$-$c$) of 512-341-10 for MaxViT-T, 1280-853-10 for EfficientNet-B0, and 2520-1680-10 for RegNetX-32GF.
+Four optimizers, Adam, NAdam, MBGD, and MBGDM, were applied to all tasks.
+Specifically, for regression tasks, we tested learning rates of ($10^{-4}$, $10^{-3}$, $10^{-2}$, $0^{-1}, 11$, 22 33).
+Larger learning rates of ($1$, 22, 33) were often excluded for the UV model due to divergence.
+For classification tasks, we tested learning rates of ($10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $0^{-1}, 11$) with Adam and NAdam, and ($10^{-3}$, $10^{-2}$, $0^{-1}, 11$, $2$, 33, 55) with MBGD and MBGDM.
+Model performances are summarized in Tables SM3, SM4, SM5 and SM6, and Figures SM17 and SM18 illustrate that UDV consistently converges to a low-rank solution.
 Generally, UDV exhibits the most pronounced decaying pattern in singular values.
@@
 Although full-batch training converges more slowly than stochastic (mini-batch) methods, it exhibits a similar singular value decay pattern, confirming our earlier observations.
-B.2 Comparison of Model Performance with the MBGDM Algorithm
-In Section 4.1, we presented the performance of each model (UDV, UV, and UV-ReLU) in terms of generalization power (measured by validation loss or validation accuracy) along with the corresponding singular value spectrum, as shown in Table 1 and Figure 3.
-The results were obtained by selecting the optimal algorithm and learning rate pair for each model.
-Here, we conduct the same experiment, but focus exclusively on the MBGDM algorithm.
-The results, presented in Table SM1 and Figure SM10 are similar to the ones in Section 4.1, showing similar trends, highlighting consistency across methods.
-Table SM1: Model performance using MBGDM. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
-Tasks Regression (Test Loss) Classification (Test Accuracy)
-Dataset HPART NYCTTD MNIST
-UDV
-$1.345\times 10^{-3}$
-MBGDM: $10^{-1}$
-$5.248\times 10^{-6}$
-MBGDM: $10^{-1}$
-M: 99.67%
-MBGDM: $10^{-2}$
-E: 99.63%
-MBGDM: $10^{-1}$
-R: 99.74%
-MBGDM: $10^{-2}$
-UV
-$1.337\times 10^{-3}$
-MBGDM: $10^{-2}$
-$5.259\times 10^{-6}$
-MBGDM: $10^{-1}$
-M: 99.69%
-MBGDM: $10^{-2}$
-E: 99.59%
-MBGDM: $10^{-1}$
-R: 99.56%
-MBGDM: $10^{-1}$
-UV-ReLU
-$1.244\times 10^{-3}$
-MBGDM: $10^{-1}$
-$5.264\times 10^{-6}$
-MBGDM: $10^{-1}$
-M: 99.63%
-MBGDM: $10^{-2}$
-E: 99.68%
-MBGDM: $10^{-1}$
-R: 99.66%
-MBGDM: $10^{-1}$
-Figure SM10: Singular value spectrum corresponding to Table SM1.
+Figure SM17: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the regression tasks.
+Figure SM18: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the MNIST dataset.
 B.3 Additional Details on the Pruning Experiment
 The SVD-based pruning experiments yield conclusions consistent with the singular value decay pattern.
-UDV typically exhibits the fastest decay, leading to more compact models while maintaining performance, as shown in Figure SM13.
-Figure SM11: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the regression tasks.
-Figure SM12: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the MNIST dataset.
-Figure SM13: The performance of SVD-based pruning.
+UDV typically exhibits the fastest decay, leading to more compact models while maintaining performance, as shown in Figure SM19.
+Figure SM19: The performance of SVD-based pruning.
 The index in x-axis represents the number of neurons in the diagonal layer after pruning.
-For the HPART and NYCTTD datasets, the validation loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$).
+For the HPART and NYCTTD datasets, the test loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$).
 Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale.
 The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice.
-For the MNIST dataset, the results show the validation accuracy after pruning the model.
+For the MNIST dataset, the results show the test accuracy after pruning the model.
 We observed that the learning rate can have some impact on the singular value decay pattern.
-In particular, a very large learning rate may cause oscillations in both training and validation loss, yet may result in a rapid decay of the spectrum.
-Conversely, a small learning rate may lead to a less pronounced spectral decay but still can yield a comparable validation loss to that of the optimal learning rates.
-For example, the difference in validation losses between the Adam optimizer with learning rates of 1e-3 and 1e-4 was negligible, yet the singular value spectra differed (see Figure SM14).
-This discrepancy also affects the performance of the pruning experiments (see Figure SM15).
-Figure SM14: Differences in singular value patterns across varying learning rates.
-Figure SM15: Different learning rates lead different pruning performance.
-the validation loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$).
+In particular, a very large learning rate may cause oscillations in both training and test loss, yet may result in a rapid decay of the spectrum.
+Conversely, a small learning rate may lead to a less pronounced spectral decay but still can yield a comparable test loss to that of the optimal learning rates.
+For example, the difference in test losses between the Adam optimizer with learning rates of $10^{-3}$ and $10^{-4}$ was negligible, yet the singular value spectra differed (see Figure SM20).
+This discrepancy also affects the performance of the pruning experiments (see Figure SM21).
+Figure SM20: Differences in singular value patterns across varying learning rates.
+Figure SM21: Different learning rates lead different pruning performance.
+the test loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$).
 Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale.
 The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice.
-When the learning rate is close to the optimal value, a ‘stair-step’ pattern (see Figure SM16) can be observed on the loss or accuracy curve in the classification task.
+When the learning rate is close to the optimal value, a ‘stair-step’ pattern (see Figure SM22) can be observed on the loss or accuracy curve in the classification task.
 It could be attributed to the model continuously searching for a low-rank solution, increasing the rank gradually.
-Figure SM16: The ‘stair-step’ accuracy curve in classification (RegNetX-32GF with Adam).
+Figure SM22: The ‘stair-step’ accuracy curve in classification (RegNetX-32GF with Adam).
 The sub-figures in the second row correspond to the respective singular value spectra, indicating that UDV consistently seeks low-rank solutions regardless of the learning rate.
@@
 We conducted a preliminary experiment to verify the feasibility of incorporating ReLU into the UDV, leaving a comprehensive study to future work.
-We derived the optimizers and learning rates from Tables SM2, SM3, SM4 and SM5, but only used the HPART dataset for regression and the MNIST for classification.
-Figure SM17 and Figure SM18 showed that UDV-ReLU and UDV-ReLU-s exhibit a similar (or even more pronounced) decaying pattern in singular values as observed in the proposed UDV.
+We derived the optimizers and learning rates from Tables SM3, SM4, SM5 and SM6, but only used the HPART dataset for regression and the MNIST for classification.
+Figure SM23 and Figure SM24 showed that UDV-ReLU and UDV-ReLU-s exhibit a similar (or even more pronounced) decaying pattern in singular values as observed in the proposed UDV.
 The ‘stair-step’ accuracy curve was also observed in these models.
-Figure SM17: ReLU in UDV: Singular value pattern on HPART dataset.
-Figure SM18: ReLU in UDV: Validation accuracy (the first row) and the singular value pattern (the second row) on MNIST dataset. Faster convergence can be achieved by fine-tuning the learning rate.
+Figure SM23: ReLU in UDV: Singular value pattern on HPART dataset.
+Figure SM24: ReLU in UDV: test accuracy (the first row) and the singular value pattern (the second row) on MNIST dataset. Faster convergence can be achieved by fine-tuning the learning rate.
 B.5 Comparison of UDV and Three-Layer Fully Connected Networks
@@
 To highlight the critical role of the UDV structure and its constraints, we include the following models in the comparison: the proposed UDV model, a model based on the UDV structure but without the constraints (UDV unconstrained), a fully connected three-layer neural network (UFV), and the standard UV model as the baseline.
-The results, also included in Tables SM2, SM3, SM4 and SM5, are shown in Figure SM19 and Figure SM20.
+The results, also included in Tables SM3, SM4, SM5 and SM6, are shown in Figure SM25 and Figure SM26.
 These findings demonstrate that the singular value decay in UDV unconstrained, UFV, and UV is significantly slower than in the proposed UDV framework.
 This underscores the critical role of the diagonal layer and constraints in facilitating low-rank solution identification within the proposed structure.
-Figure SM19: Comparison of singular value spectrum among UDV, UDV (unconstrained),UFV and UV on the regression tasks.
-Figure SM20: Comparison of singular value spectrum among UDV, UDV (unconstrained), UFV and UV on the classification datasets (MNIST).
+Figure SM25: Comparison of singular value spectrum among UDV, UDV (unconstrained),UFV and UV on the regression tasks.
+Figure SM26: Comparison of singular value spectrum among UDV, UDV (unconstrained), UFV and UV on the classification datasets (MNIST).
 B.6 Comparing the Effects of UDV and Weight Decay Regularization
@@
 For the regularization parameter ($\gamma$), we tested values $10^{-6}$, $10^{-5}$, $\dots$, $10^{-1}$.
-Increasing $\gamma$ led to a faster decay in the singular value spectrum but at the cost of higher training and validation losses.
-Figure SM21 illustrates an example where we selected the largest $\gamma$ values for which the validation loss remained within 10% of the UDV baseline.
+Increasing $\gamma$ led to a faster decay in the singular value spectrum but at the cost of higher training and test losses.
+Figure SM27 illustrates an example where we selected the largest $\gamma$ values for which the test loss remained within 10% of the UDV baseline.
 The resulting singular value decay in the UV and UFV models was less pronounced than in the UDV model.
-Further increasing $\gamma$ to match or exceed the spectral decay of the UDV model resulted in significantly worse validation loss compared to the UDV network.
-The complete results are presented in Figures SM22 and SM23.
-Figure SM21: Comparison of singular value spectrum among the proposed UDV, UFV and UV (with weight decay) on the HPART dataset.
-In addition to the UDV baseline, the UFV and UV models include the regularization parameter $\gamma$ and the relative change in validation performance.
+Further increasing $\gamma$ to match or exceed the spectral decay of the UDV model resulted in significantly worse test loss compared to the UDV network.
+The complete results are presented in Figures SM28 and SM29.
+Figure SM27: Comparison of singular value spectrum among the proposed UDV, UFV and UV (with weight decay) on the HPART dataset.
+In addition to the UDV baseline, the UFV and UV models include the regularization parameter $\gamma$ and the relative change in test performance.
 The relative change is defined as ($\frac{loss_{model}-loss_{baseline}}{loss_{baseline}}\times 100\%$), representing how much worse the model performs compared to the UDV baseline.
-Figure SM22: Extension of Figure SM21.
+Figure SM28: Extension of Figure SM27.
 Comparison of singular value spectrum between the proposed UDV and UFV (with weight decay) on the HPART dataset.
 Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
-Figure SM23: Extension of Figure SM21.
+Figure SM29: Extension of Figure SM27.
 Comparison of singular value spectrum between the proposed UDV and UV (with weight decay) on the HPART dataset.
 Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
-Table SM2: Experiments using Adam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
+B.7 A Toy Example of Causal Language Modeling Using a Pre-trained LLaMA-2 Model with LoRA-Based UDV Fine-Tuning
+To further broaden the application scope of the proposed method, we integrate UDV into the pre-trained model using the LoRA framework and evaluate its effectiveness on the causal language modeling task (i.e., a generative task where a model predicts the next token in a sequence based solely on past tokens).
+Experiments are conducted using the LLaMA-2-7B model (Touvron et al., 2023) and the WikiText-2 dataset (Merity et al., 2016), which is a common benchmark test.
+LoRA (Low-Rank Adaptation) is a parameter-efficient fine-tuning technique that inserts trainable low-rank matrices into pre-trained models (Hu et al., 2022).
+By freezing the original model weights and only training a small number of additional parameters, LoRA enables efficient adaptation to downstream tasks while significantly reducing computational and storage costs.
+Specifically, LoRA introduces trainable low-rank matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ as a side path to the frozen pre-trained weight matrix $W_{0}$, and expresses the forward pass as $h=W_{0}x+BAx$, where $x$ is the input.
+Here we consider $d=k$ for simplicity.
+By inserting a diagonal matrix $D$ between the matrices $B$ and $A$, the LoRA structure naturally extends to the LoRA-based UDV structure, as illustrated in Figure SM30.
+After the weight update of $U,D$ and $V$, constraints described in Equation 5 are applied to the LoRA-based UDV structure.
+In practice, the diagonal matrix can be efficiently implemented using a vector $d\in\mathbb{R}^{1\times r}$, further reducing computational and memory overhead.
+Consequently, the LoRA-based UDV introduces a negligible increase in the number of trainable parameters compared to the original LoRA structure, as it adds only $r$ parameters per LoRA layer replacement.
+Figure SM30: LoRA forward pass (Left) and LoRA-based UDV forward pass (Right).
+The inputs and pre-trained weights are identical in both cases.
+Matrices with the same fill color indicate matching shapes.
+We initialized both the LoRA and LoRA-based UDV structures from the same starting point: $B=V=0$, $A=U=\mathcal{N}(0,0.01)$ and $D=I$, where $I$ is the identity matrix.
+Mini-batch gradient descent (MBGD) with learning rates $lr=0.001$, using an effective batch size of 32, was applied to better observe the effects of the imposed constraints.
+Following the standard LoRA setup, only the linear layers associated with the Query ($Q$) and Value ($V$) projections in the attention mechanism were replaced with LoRA and LoRA-based UDV structures.
+We evaluated the models using rank values $r\in\{4,8,16,32,64\}$ , where the dimensionality of $d$ depends on the shape of the original layer.
+Both structures were fine-tuned for 30 epochs, and then pruning was applied as described in Section 3.1.3.
+Instead of applying uniform pruning across all LoRA-based UDV structures, we set several thresholds on the retained energy—defined as the sum of the squared singular values—to dynamically prune each UDV block based on its individual importance.
+Note that the pruning can be applied to the LoRA structure in the same manner as the LoRA-based UDV structure by temporarily inserting an frozen identity matrix as a diagonal layer.
+Perplexity (i.e., the exponential of the average negative log-likelihood) is used as the primary evaluation metric for this experiment, and lower perplexity indicates better language modeling performance.
+Furthermore, the pruning performance was demonstrated by the perplexity change in terms of the compression ratio (i.e., the ratio between the number of trainable parameters of the compressed structure and the original structure)
+Since the hard constraints slower the convergence, and its fine-tuning performance is slightly lower than that of the original LoRA structure within the same number of epochs, as shown in Figure SM31 shown.
+However, the pruning results (see Figure SM32) demonstrate that the LoRA-based UDV is significantly more robust to high compression ratios.
+Even at a compression ratio of $0.1$, the LoRA-based UDV is able to maintain performance effectively.
+Figure SM31: Comparison of perplexity between LoRA and LoRA-based UDV during training. The perplexity at epoch 0 reflects the performance of the pre-trained model before any fine-tuning.
+Figure SM32: Comparison of pruning performance between LoRA and LoRA-based UDV during training.
+This toy example is not intended to present a mature pruning method, but rather to highlight the potential of the proposed UDV block, which can be applied to a wide range of models and demonstrates promising pruning performance.
+Appendix C Extended Discussion on the Related Work
+Burer-Monteiro factorization.
+BM factorization was proposed for solving semidefinite programs (Burer & Monteiro, 2003, 2005; Boumal et al., 2016, 2020; Cifuentes, 2021) and has been recognized for its efficiency in addressing low-rank optimization problems, e.g., (Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017, 2018; Hsieh et al., 2018; Sahin et al., 2019; Lee et al., 2022; Yalçın et al., 2023).
+Building on the connections between training problems for (non-linear) two-layer neural networks and convex (copositive) formulations (see (Pilanci & Ergen, 2020; Ergen & Pilanci, 2020; Sahiner et al., 2021) and the references therein), Sahiner et al.
+(2023) recently introduced BM factorization to solve convex formulations of various architectures, including fully connected networks with ReLU and gated ReLU (Fiat et al., 2019) activations, two-layer convnets (LeCun et al., 1989a), and self-attention mechanisms based on Transformers (Vaswani et al., 2017).
+Implicit regularization.
+One promising line of research that aims to explain the successful generalization abilities of neural networks follows the concept of ‘implicit regularization,’ which results from the optimization methods and formulations used during training (Neyshabur et al., 2014, 2017; Neyshabur, 2017).
+Several studies explore matrix factorization to investigate implicit bias (Gunasekar et al., 2017; Arora et al., 2018; Razin & Cohen, 2020; Belabbas, 2020; Li et al., 2021).
+Le & Jegelka (2022) extend these results to the final linear layers of nonlinear ReLU-activated feedforward networks with fully connected layers and skip connections.
+More recently Timor et al.
+(2023) investigated implicit regularization in ReLU networks.
+Much of the existing work focuses on gradient flow dynamics in the limit of infinitesimal learning rates.
+In particular, Gidel et al.
+(2019) examined discrete gradient dynamics in two-layer linear neural networks, showing that the dynamics progressively learn solutions of reduced-rank regression with a gradually increasing rank.
+Constrained neural networks.
+Regularizers are frequently used in neural network training to prevent overfitting and improve generalization (Goodfellow et al., 2016), or to achieve structural benefits such as sparse and compact network architectures (Scardapane et al., 2017).
+However, it is conventional to apply these regularizers as penalty functions in the objective rather than as hard constraints, addressing them within gradient-based optimization via their (sub)gradients.
+This approach is likely favored due to the ease of implementation, as pre-built functions are readily available in common neural network packages.
+Several independent studies have also applied proximal methods across different networks and applications, finding that proximal-based methods tend to yield solutions with more pronounced structures (Bai et al., 2019; Yang et al., 2020; Yun et al., 2021; Yurtsever et al., 2021; Yang et al., 2022).
+Structural regularization in the form of hard constraints, however, appears to be rare in neural network training.
+One notable exception is in the context of neural network training with the Frank-Wolfe algorithm (Pokutta et al., 2020; Zimmer et al., 2022; Macdonald et al., 2022).
+Recently, Pethick et al.
+(2025) revealed parallels between Frank-Wolfe on constrained networks and algorithms that post-process update steps, such as Muon (Jordan et al., 2024), which achieves state-of-the-art results on nanoGPT by orthogonalizing the update directions before applying them.
+There are other examples of constraints in neural networks.
+For instance, constraints can be applied to ensure adherence to real-world conditions in applications where they are necessary (Pathak et al., 2015; Márquez-Neila et al., 2017; Jia et al., 2017; Kervadec et al., 2019; Weber et al., 2021), or to incorporate physical laws in physics-informed neural networks (Raissi et al., 2019; Lu et al., 2021; Patel et al., 2022).
+In these cases, the feasible set is typically complex and difficult to project onto; therefore, optimization algorithms like primal-dual methods or augmented Lagrangian techniques are used.
+Constraints are also present in lifted neural networks, a framework where the training problem is reformulated in a higher-dimensional ‘lifted’ space.
+In this space, conventional activation functions like ReLU, used in the original formulation, are expressed as constraints (Askari et al., 2018; Sahiner et al., 2021; Bartan & Pilanci, 2021; Prakhya et al., 2024).
+Pruning.
+Neural networks are overparameterized, which can enhance generalization and avoid poor local minima.
+But such models then suffer from excessive memory and computational demands, making them less efficient for deployment in real-world applications (Chang et al., 2021).
+Several compression techniques have been studied in the literature, including parameter quantization (Krishnamoorthi, 2018), knowledge distillation (Gou et al., 2021), and pruning.
+Pruning reduces the number of parameters, resulting in more compact and efficient models that are easier to deploy.
+The literature on pruning is extensive, with diverse methods proposed with distinct characteristics.
+Various criteria are used to determine which weights to prune, including second-order derivative-based methods (LeCun et al., 1989b; Hassibi & Stork, 1992), magnitude-based pruning (Janowsky, 1989; Han et al., 2015), saliency heuristics (Mozer & Smolensky, 1988; Lee et al., 2018), and matrix or tensor factorization-based techniques (Xue et al., 2013; Sainath et al., 2013; Jaderberg et al., 2014; Lebedev et al., 2015; Swaminathan et al., 2020), among others.
+Pruning by singular value thresholding has shown promising results, particularly in natural language processing (Chen et al., 2021), and is often used along with various enhancements such as importance weights and data whitening for effective compression of large language models (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024).
+A comprehensive review of pruning methods is beyond the scope of this paper due to space limitations and the diversity of approaches.
+For more detailed reviews, we refer to (Reed, 1993; Blalock et al., 2020; Cheng et al., 2024), and the references therein.
+Table SM3: Experiments using Adam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
 Tasks Regression (Test Loss) Classification (Test Accuracy)
@@
 HPART
-(×$\times 10^{-3}$)
+($\times 10^{-3}$)
 NYCTTD
-(×$\times 10^{-6}$)
+($\times 10^{-6}$)
 MNIST
-(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
-(×$\times 100\%$)
+(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
+($\times 100\%$)
 LR: $10^{-6}$
@@
 LR: $5\times 10^{0}$
-Table SM3: Experiments using NAdam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
+Table SM4: Experiments using NAdam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
 Tasks Regression (Test Loss) Classification (Test Accuracy)
@@
 HPART
-(×$\times 10^{-3}$)
+($\times 10^{-3}$)
 NYCTTD
-(×$\times 10^{-6}$)
+($\times 10^{-6}$)
 MNIST
-(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
-(×$\times 100\%$)
+(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
+($\times 100\%$)
 LR: $10^{-6}$
@@
 LR: $5\times 10^{0}$
-Table SM4: Experiments using MBGD optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
+Table SM5: Experiments using MBGD optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
 Tasks Regression (Test Loss) Classification (Test Accuracy)
@@
 HPART
-(×$\times 10^{-3}$)
+($\times 10^{-3}$)
 NYCTTD
-(×$\times 10^{-6}$)
+($\times 10^{-6}$)
 MNIST
-(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
-(×$\times 100\%$)
+(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
+($\times 100\%$)
 LR: $10^{-6}$
@@
 R: 99.55
-Table SM5: Experiments using MBGDM optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
+Table SM6: Experiments using MBGDM optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
 Tasks Regression (Test Loss) Classification (Test Accuracy)
@@
 HPART
-(×$\times 10^{-3}$)
+($\times 10^{-3}$)
 NYCTTD
-(×$\times 10^{-6}$)
+($\times 10^{-6}$)
 MNIST
-(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
-(×$\times 100\%$)
+(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
+($\times 100\%$)
 LR: $10^{-6}$